### v1

# Tadashi: Enabling AI-Based Automated Code Generation With Guaranteed Correctness

Abstract.
Frameworks and DSLs auto-generating code have traditionally relied on human experts developing them to have in place rigorous methods to assure the legality of the applied code transformations. Machine Learning (ML) is gaining wider adoption as a means to auto-generate code optimised for the hardware target. However, ML solutions, and in particular black-box DNNs, provide no such guarantees on legality. In this paper we propose a library, Tadashi, which leverages the polyhedral model to empower researchers seeking to curate datasets crucial for applying ML in code-generation. Tadashi provides the ability to reliably and practically check the legality of candidate transformations on polyhedral schedules applied on a baseline reference code. We provide a proof that our library guarantees the legality of generated transformations, and demonstrate its lightweight practical cost. Tadashi is available at https://github.com/vatai/tadashi/.
Tadashi sampling the space of transformations
1. Introduction
Machine learning (ML) is disrupting all forms of digital media, and the creation of computer programs is no exception. ML can be and has been applied in various ways with the aim of making the work of programmers (Dehaerne et al., 2022a) easier: generating code from “natural languages” (Phan, 2019), write a summary (e.g., write commit messages) of a piece of code (Shido et al., 2019), auto-complete code on-the-fly while it’s being written (Github, [n. d.]) etc. In this paper, we consider the task of optimising code to a hardware target, i.e., using ML to transform the hotspots in the code, i.e., loops, in a way that improves performance.
Checking if the program generated by the ML model is faster then the baseline is just the matter of running it, then comparing against the run-time of the original code. However, it is a major challenge to get formal guarantees that the transformed code produces the same output as the original (Dehaerne et al., 2022b; Munoz et al., 2024; Jiang et al., 2024). This impedes the adoption of ML for code generation and transformation in mission critical scientific/engineering codes, or codes that come with high liability, such as in the financial, legal, and corporate sectors. Additionally, there is a lack of tooling for generating and sharing parameterised datasets comprised of legal auto-generated code, which further constraints progress in the field since, to date, ML-based solutions have to curate their datasets.
To this end, we developed Tadashi, a Python library for sampling the search space of possible loops transformations, with the capability to verify the legality (i.e., correctness) of the sampled transformations. Tadashi allows the user to programmatically specify the pattern for sampling legal individual or combinations of transformations, based on the preference of the ML method that would be used to explore the search space of transformations. Through the use of the polyhedral model, Tadashi is capable of enabling ML engineers to explore novel and more aggressive methods of code optimisation in the datasets they curate, while guaranteeing that the results of the generated code are functionally identical to the original program. Tadashi key point is the ease-of-use for ML practitioners; Tadashi aims to provide a simple Python interface, which can be used to curate datasets or perform direct code-generation to be used with different ML approaches.
In this paper, we make the following contributions:
We introduce Tadashi, a Python library designed for ML practitioners seeking to design tools for code auto generation with correctness guarantees. Tadashi lowers the barrier to entry of polyhedral loop transformations; a model that can give mathematically provable guarantees on the correctness (i.e., legality) of the transformed code.
We demonstrate how Tadashi can be used by an ML system by implementing an end-to-end stub reinforcement learning system.
We evaluate performance to confirm it is not an obstacle in Tadashi’s adoption. We observe throughput measurements above 32 transformations and legality checks per second, going up to 781 for some benchmarks. The overheads of parsing and generating code is negligible in comparison to compiling and running the generated code.
2. Background and Motivation
In this section we discuss the motivation for designing a library that leverages the polyhedral model as bases for guaranteeing correctness of ML-driven code generation.
2.1. Correctness in ML-Based Code Auto-Generation
ML is being increasingly used in code auto-generation, at different levels of abstraction, e.g., source code (Stein et al., 2020), abstract syntax tree (AST) (Shido et al., 2019), polyhedral (Baghdadi et al., 2021), dependency graphs (Cummins et al., 2020), intermediate Representations (IR) (Ben-Nun et al., 2018), and assembly instructions (Mankowitz et al., 2023).
For guarantees on correctness, existing approaches for using ML in code generation or transformation rely on one of the following approaches:
Extensive unit testing: for this, widely used, approach, a suite of unit tests is used to validate correctness (Cummins et al., 2024; Beer et al., 2024; Du et al., 2024; Chen et al., 2021). While unit testing can be effective, it provides no guarantees on correctness111”Program testing can be used to show the presence of bugs, but never to show their absence!”– Edsger W. Dijkstra.
In summary, the approaches above either provide no formal guarantee on functional correctness, or confine the user to a very constrained form of code generation. In addition, we observed that for the above approaches, it was the burden of the researchers to implement a sampling strategy and curate the dataset. Solving those two problems, correctness and agency, is the motivation for Tadashi.
2.2. The Polyhedral Model
Polyhedral techniques have a long history with one of the seminal papers dating back to 1996 (Feautrier, 1996) and are well established and used in compilers (Aho et al., 2007). The polyhedral model is a very efficient way to manipulate and reason about nested loops, including the capability to check the legality222Legality is the technical term, used in the polyhedral literature with a precise definition which translates to the transformed code calculating the same results as the original, i.e., functional correctness. of the transformations. It has been used in optimization pass of compilers (Grosser et al., 2012), as well as standalone source-to-source optimization tools (Bondhugula et al., 2008; Verdoolaege et al., 2013) relying on integer linear programming with some optimization heuristic.
The polyhedral model is a representation for loop nests with regular boundaries and memory accesses, know as static control parts (see Listing 1 an example and Section 2.3 for a description). The polyhedral representation separates the statement instances from the order in which they are executed: Each of these instances is then mapped to an integer tuple, a timestamp, and then the statement instances are ordered according the lexicographical order of their timestamps. For a given statement , the set of statement instances is called the domain and is denoted as ; while the mapping which determines their order is the schedule, which is denoted by . Listing 1 and Figure 2(d) give a simple example of a SCoP and possible schedules.
The main idea of polyhedral optimization is to transform the schedule (or create a new one from scratch) which reorders the statement instances so they are executed more efficiently on a given hardware (e.g., by increasing memory locality). As one could expect, not all schedule transformations are legal, however checking legality is very convenient in the polyhedral model. Dependency analysis (Feautrier, 1991) yields a directed dependecy graph . The vertices of are statements (or equivalently their domains), and the edges are labelled with maps which describe the dependencies on the level of statement instances such as
Applying the schedule to the dependee and dependent of these maps takes them into the time domain, and (1) becomes ( and are timestamps). If lexicographically precedes , i.e.,
then the schedule is legal, because it maps the dependent to a later time then the timestep of the dependee. Because (2) is equivalent to , we can (symbolically) compute and if is lexicographically positive, the schedule is legal. Computing is useful because it provides us with even more useful information: if an entry in is zero, it means there are no dependencies along the corresponding dimension of the schedule, and consequently the generated loop for that dimension can be parallelised.
Example schedules
2.3. Constraints of the Polyhedral Model
To be able to represent the domain, schedule, and dependencies efficiently, the source codes represented by the polyhedral model are restricted to Static Control Parts, or SCoPs for short (Listing 1 is a simple example). While these restrictions are not negligible, most of high performance and science codes do not violate them. The main restrictions of SCoPs are the following: (1) The control-flow must be static, which means it does not depend on the input, other than parameters (such as loop boundaries) which remain constant during execution, (2) there is no memory aliasing, no pointer manipulation, and (3) the symbolic expressions and formulas, used throughout the model, need to be expressed using a matrix and a vector as a (quasi-)affine expression, i.e., (sometimes referred to as Presburger formulas (Verdoolaege et al., [n. d.])). For example, the domains are expressed as ; schedules are (quasi-)affine functions . The polyhedral model can also handle tiling, so the formulas include integer division with a constant as well, and this is what the “quasi-” prefix refers to. However, for brevity and clarity, we omit discussion of these details and for a more comprehensive description of the polyhedral model we recommend (Feautrier, 1996, 1997; Verdoolaege et al., [n. d.]).
Abiding by those restrictions, which is reasonably possible in a wide range of high performance science and engineering applications, enables ML systems that use Tadashi to guarantee correct and legal output at a scale, unlike most other existing solutions.
3. Tadashi: A Library to Sample Legal Polyhedral Transformations
Tadashi is a library that provide the means for ML practitioners to sample the space of legal polyhedral loop transformations. The user can use Tadashi to describe a sampling strategy that is suited for the ML approach, e.g. uniform sampling to curate datasets vs. guided exploration of the space of legal transformations. Figure 1 gives an overview of Tadashi. At an architectural level, Tadashi consists of a Python front-end and a C/C++ back-end (aptly named ctadashi). The front-end provides a minimal and easy-to-use interface to sample from the space of possible transformations, check their legality, and generate the transformed code. More importantly, the front-end being in Python encourages fast adoption by ML practitioners since Python is the de facto language used in AI/ML frameworks. The back-end implements this functionality, using the Integer Set Library (ISL) (Verdoolaege, 2010) and Polyhedral Extract Tool (PET) (Verdoolaege and Grosser, 2012) polyhedral libraries.
Existing ISL wrapper libraries for Python expose the full API of the library, which can be daunting for users not versed in the concepts of polyhedral compilation. The goal of Tadashi is to lower this barrier to entry for ML engineers and researchers by providing out-of-the-box implementations of transformations, legality check and code generation. Moreover, the functionality is exposed in a more idiomatically Pythonic way. For example, a tree of nested loops is presented is a hierarchy of nested objects similar to how HTML parse trees are used in Python libraries, i.e., a transformation on a given node can be done by calling an appropriate method of a given object, while hiding all bookkeeping necessary for interacting with ISL and PET libraries.
3.1. Python Front-End: Polyhedral Made Easy
To demonstrate the ease of use of Tadashi, Listing 2 performs a tiling transformation, checks legality, and generates new code in less than 10 lines of a fully functional Python example. Line 2 creates the app object, which stores the information required by Tadashi to load the input file. The scops object in line 3 stores the polyhedral representations of the loop nests. The node object on line 4 specifies the loop inside the first SCoP, and a tiling transformation (with tile size 16) is performed on line 5. The legality of the tiling is checked together with the transformation, and the result is stored in the legal variable. If the transformation is legal, the new code is generated in line 7.
To pass an input file to Tadashi, it must be stored in an instance of the App abstract base class. Currently Tadashi provides two implementations of App: the Simple class (used in line 2) which can be used for simple single-file programs and Polybench (used later in Section 5) to load Polybench apps (Reisinger, 2024). The classes derived from App store the input filename to Tadashi and additionally include directories are also needed by Tadashi when parsing the file. For convenience, App also defines compile() and measure() methods.
The polyhedral representation of all loop nests (SCoPs) are implemented as an array-like scops object (line 3), and from the first SCoP (scops[0]) a node of the schedule trees (Verdoolaege et al., 2014) is selected (line 4). For a user not familiar with the polyhedral model and schedule trees, a node is “an entity that can be transformed”, which usually means a loop in the loop nest (a discussion of schedule trees can be found in Section 4). Since some transformations are not available for certain loops, the Node class provides valid() and valid_args() methods to check availability. The nodes of a scop can be accessed through the (dynamic) scop.schedule_tree[] as a list, or as a tree using the goto_child(child_idx) and goto_parent() methods of nodes.333The root is scop.schedule_tree[0]!.
A code transformation is performed by passing a TrEnum (an Enum type enumerating all transformations) object to node.transform() (with the transformation-specific argument). A list of transformations with a short description of their arguments can be found in Table 1.
These are primitive transformations, in the sense that they are composable, so we can apply a sequence of transformations on the same SCoP. In the current version of Tadashi, we prioritised the the implementation of loop transformations that have high impact on performance (Pouchet et al., 2011)
The return value of the transform() method is the result of the legality check and, based on it, the users can decide what they want to do. In case of a legal transformation, one likely action would be generating the C code (line 7), compiling it and measuring the execution time, while for an illegal transformation one could invoke the rollback() method of the node, to return the polyhedral representation to legal state.
TrEnum |
Args |
TILE |
tile size |
INTERCHANGE |
FULL_FUSE |
FUSE |
2 loop indices |
FULL_SHIFT_VAL |
const shift value |
PARTIAL_SHIFT_VAL |
stmt idx, const shift value |
FULL_SHIFT_VAR |
coeff, shift iter idx |
PARTIAL_SHIFT_VAR |
stmt idx, coeff, shift iter idx |
FULL_SHIFT_PARAM |
coeff, shift param idx |
PARTIAL_SHIFT_PARAM |
stmt idx, coeff, shift param idx |
SCALE |
coeff |
SET_PARALLEL |
SET_LOOP_OPT |
AstLoopType enum |
The Python front-end, aims to hide the details of the polyhedral model. To this end, Tadashi provides aliases to class and variable names which might be daunting or confusing to a user unfamiliar with polyhedral concepts (see Table 2). This way, Tadashi can hide the details of what is happening under the hood, which we describe in Section 4.
3.1.1. Sampling Strategies
Listing 2 demonstrates the application of a single primitive transformation. However, in consequence of the compositionality of polyhedral transformations, the user can apply a sequence of transformations and legality checks to a SCoP (before generating the code) and hence obtain a different composite transformation. This is a new sample from the space of polyhedral transformations, different from its component primitive transformations. Put differently, sampling the space of polyhedral transformations with Tadashi consists of applying a sequence of transformations to a SCoP, where each entry consists of specifying a location (a node in the schedule tree), the transformation (TrEnum) and a set of arguments. A location can be selected using the schedule_tree[] member or with a combination of goto_* methods of the Node class. The transformation can be selected from the possible values TrEnum values and checked if it is available on the selected node. Finally, selecting an argument is somewhat convoluted, since the available values depend not only on the selected transformation, but also on the concrete properties of the selected schedule node (such as its partial schedule, type and number of children etc.). Tadashi exposes the necessary information for random argument selection by providing the list of available values or the lower and upper bounds for integer arguments.
For example, if we intend to narrow the search space by iterating over combinations of shifts and tiles, shifting by single increments of surrounding iteration variables and tile sizes of powers of starting from to , the following steps need to be taken. The desired loop is saved as a node, and each transformation is combined with its arguments. For tiling, this can easily be accomplished using list comprehension, by gathering (TrEnum.TILE, 2**i) pairs where i iterates over range(3, 11). To create the argument list of the shift transformation, the user obtains the lower and upper bounds of all arguments for FULL_SHIFT_VAR on node. The first argument, the coefficient, is set to since we want single increments but the second argument, the iterator index is varied, i.e., list of triplets is created containing FULL_SHIFT_VAR, 1 and iter_idx, where iter_idx iterater over the range between the lower and upper bound of the second argument. Using the product() function from Python’s itertools package, these two lists are combined into a list of (tile, shift) pairs. Then by iterating through these pairs in a loop, we can call the node’s transform() method, first with *tile, then with *shift, followed by a checking the legality and generating code.
| Polyhedral | Simplified |
Scops |
LoopNests |
Scop |
LoopNest |
scop.schedule_tree[] |
loop_nest.nodes[] |
3.2. Using Different ML Methods With Tadashi
The goal of Tadashi is to lower the barrier to entry for user-guided sampling of legal loop transformations. Tadashi does not in itself implement any type of ML; we do however design Tadashi with considerations to how various ML approaches can use it. The design of Tadashi supports the following ML search space exploration approaches:
3.2.1. Reinforcement Learning (RL)
RL (Sutton and Barto, 2018) is a prime candidate to use Tadashi, since Tadashi’s interface maps naturally to RL concepts. Similar to choosing winning moves in a game, an RL agent chooses promising transformations. The transformed code is executed by the runtime and speedup over the original version serves a the reward signal. Sequences of transformations can be sampled by using approaches such as Monte Carlo tree search (MCTS), where a neural network model gives scores to possible choices - an approach similar to Alpha Go(Silver et al., 2016) system.
3.2.2. Supervised Learning
Different versions of transformed source codes and corresponding speedups can be saved in a database. A model to predict speedups from transformations can be trained in supervised fashion to be used as a value model in an off-policy reinforcement learning setup or in other applications.
Such approach is especially valuable, because evaluation of transformed code takes significantly more time, compared to selecting and performing transformations (see sec. 5)
The choice of transformations can be done in random or by grid search. Many transformations require additional corresponding parameters, e.g. choosing to tile a loop requires choosing a tile size. These arguments (second column in Table 1) allow the user to combinatorially increase the coverage of the search space to consequently increase the coverage of the curated dataset. For example, starting with a limited set of input C source files containing loop nests, one could run Tadashi to select random transformations (on a random input file, on random nodes of the schedule tree, with random parameters), then store the new code with the result of the legality check and the measured execution time. The dataset generated this way would consist of (input, output) pairs, where the input is (the name of) a C source file, and the output consists of: (1) the transformation, (2) transformed code, (3) legality of the transformation, (4) execution time of the transformed program.
In the case of legal transformations, the transformed code can be added to the pool of input files, expanding the number of different input files ad infinitum.
Another, approach is to sample sequences of transformations, i.e., generate a sequence (of a random length) of random transformations, where the individual transformations are sampled similarly as described in the previous paragraph. The legality check, code generation and performance measurement are performed only at the end of the sequence.
With basic Python programming, e.g., using the itertools package, one can easily generate a regular “grid” of transformations, which is adequate for scenarios similar to auto-tuning.
3.2.3. Evolutionary Algorithms
Similar to RL, evolutionary algorithms use a mix of exploration and exploitation to traverse the search space, however without relying on on learnable parameters and performing search using more direct heuristics (Cai et al., 2023). Exploration refers to searching for new, diverse solutions by exploring a wide range of the solution space, while exploitation focuses on refining and improving existing solutions by intensively searching around the best solutions found so far. As mentioned above in other ML approaches, Tadashi’s sampling of the search space can be parameterised by the user to explore a wide range of the search space (exploration) or intensive search in a narrow region of the search space (exploitation). The runtime of the transformed code variants would be used as the objective function.
The exploration capability of evolutionary algorithms can in some instances improve if infeasible solutions appear in the population (While and Hingston, 2013), and then later be penalised to disappear over time from the population (Coello, 2022). Tadashi internally manipulates the schedule to generate both legal and illegal transformations, where we then filter out the illegal transformations. Since illegal transformations can be of value for the exploration capability of evolutionary algorithms, we allow the user to sample illegal transformations simply by not rolling them back and generating code disregarding the result of the legality check.
3.2.4. With LLM Agents
The most performant transformations identified by any of the aforementioned methods can be used as training data for Large Language Models for better end-to-end code generation or as additional signal to improve overall capacity of the model (Du et al., 2024).
However, in line with the motivation for this paper, an even more promising application of Tadashi would be to be called directly by an LLM agent in a “tool calling” paradigm (Schick et al., 2023), similar to Retrieval Augmented Generation (RAG) systems (Lewis et al., 2020) or Chain of Thought approach (Wei et al., 2024).
3.2.5. Auto-Tuning
For practical reasons, ML approaches used in auto-tuning (Mustafa, 2022; Ashouri et al., 2018) mostly rely on local search methods with focused search around particular regions in the search space, after aggressively pruning the search space. Similar to the exploitation in evolutionary algorithms, Tadashi can be setup to narrow down the search around the region for auto-tuning, based on the post-pruning range of interest.
3.2.6. ML-driven deployment of applications on new hardware
In addition to supporting the usage patterns of the ML approaches listed above, Tadashi can drive an ML approach to learnhardware specific optimisations on new hardware. After the training phase, applications ported to the new hardware can do inference on the ML model, leveraging Tadashi, to optimise the applications for the new hardware.
3.3. Extensibility of Tadashi
Tadashi is designed to be extensible w.r.t. the supported transformations, programming languages, and target hardware.
We implemented the set of transformations listed in Table 1. To achieve a larger coverage, further transformations (e.g. loop fission) could be added. This requires implementing a function which takes a schedule node as a parameter, performs the desired transformation and returns the transformed node (for details see Section 4.4.2). There are also features of ISL we did not explore yet, such as, expansion and extension nodes. A promising direction is to generate arbitrary partial schedules (which can be virtually any quasi-affine map444Namely isl_multi_union_pw_aff!.) for the band nodes in the schedule tree, providing access to space of all polyhedral transformations.
In essence, Tadashi’s core capabilities involve operating on ISL schedules, and it is not of importance what tool was used to generate them (the frontend) or how the output code is generated (the code generation backend). This means that while we currently use PET and ISL for those stages, which enables us to target C code, in the future we can integrate with compiler solutions such as Polly (Grosser et al., 2012) and Polygeist (Moses et al., 2021) to tap into their respective LLVM and MLIR ecosystems, or other tools which use ISL schedules as an intermediate representation (Vasilache et al., 2018; Verdoolaege et al., 2013) to gain additional capabilities w.r.t input languages or target hardware.
One potential scalability issue in utilising Tadashi for generating a large amount of runtime measurements is the availability of a wide range of inputs to run the generated code with. Benchmarks usually contain a small set of inputs for running the application, however, needing to run the entire application in order to time only specific small portions of it (scops) is wasteful. Since we already isolate these points of interest (scops), existing frameworks for automatic input generation (Ivanov et al., 2024; Kind et al., 2022) can be utilised to synthesise inputs which can be used to execute and time only the code in the scop.
4. Implementation of Tadashi
We now revisit the end-to-end example of Listing 2, but with our attention directed to implementation details and what is happening under the hood, in the ctadashi back-end depicted in Figure 3.
Overview of the ctadashi back-end.
4.1. The App Class
The App class (Listing 2 line 2) is completely written in Python and provided mostly for convenience. It’s main functions are to provide the source_path and include_paths properties, the former being the C source file which will be transformed, and the later a list of paths used to set the required environment variable,555Namely C_INCLUDE_PATH!. which is necessary, because when PET runs a compiler pass on the input file, it will throw an error if it cannot find an included header file.
4.2. SCoPs
The creation of the Scops object (line 3) sets things in motion in the C/C++ back-end of Tadashi. The input file is passed through PET’s C source transformations function666Namely pet_transform_C_source!. which detects all the SCoPs in the file, but instead of doing any transformations, the pet_scop objects passed to the callback are used to populate the global SCOP_INFO vector. Each entry in this vector is a scop_info_t struct, and stores the pointer to the pet_scop, the dependencies computed by ISL according to the instruction found in (Verdoolaege et al., [n. d.]) and using the stricter “may” dependencies,777Using isl_union_flow_get_may_dependence!. and pointer to the “current node”888Of type isl_schedule_node! in the schedule tree.
4.3. Schedule Tree and Navigation
Transformations in Tadashi are performed on schedule trees, which are alternative ways to represent the domains and schedules in a SCoP (see Section 2.2). The leaves of the schedule tree are the statements, the root is the union of their domains. The loops surrounding a statement (in the source code) correspond to band nodes in the tree. On the path from the root to the leaf, each band node has a partial schedule and the complete schedule of the statement (i.e., leaf) can be constructed by concatenating the partial schedules. Other nodes in the schedule tree are sequence and filter nodes. Each sequence node has multiple filter nodes as children and together they represent a sequence of statements and/or nested loops. The sequence node is the placeholder node for the entire sequence, while its children act as roots, i.e., domains for the sub-schedule trees.
Each transformation in Tadashi is defined on a node of a schedule tree, so line 4 in Listing 2 involves the following mechanism. First, scops[i] returns a Python Scop object which stores the index of the corresponding in SCOP_INFO entry on the C++ side. This Scop object wraps all the C++ functions exposed to Python by calling them with a scop_idx parameter from an object method with same name. These functions include navigation operations999Namely goto_root()!, goto_parent()!, and goto_child(idx)!. to update the current node to point to its parent, its child, or to the root of the schedule tree (for the selected scop_info_t object). The schedule_tree property on line 4 is dynamically generated, and exposes the schedule tree as a list of nodes by doing a depth-first traversal. During the traversal, we build a list of child indexes which enables Tadashi to locate any node by calling goto_child for each element in the list (implemented as the locate method of a Node). In addition to the navigation methods, Tadashi exposes wrappers to query properties of nodes, (such as node type, number of children, partial schedule, parameters, enclosing iterator variables etc.), and so enables the Python Node object to emulate the ISL schedule tree and its nodes in C++.
4.4. Transformations
After updating the current node to point to the desired location, invoking node.transform() calls the C++ function implementing the corresponding transformation. The transformation is performed on the current node of the SCoP specified by the index obtained from the node (similar to the navigation and query functions from the previous paragraph). To be able to rollback the transformation, the current node is copied, and restored if the user calls the rollback method (usually after the transformation turns out to be illegal).
4.4.1. Implementation Details of Individual Transformations
Multi dimensional tiling is already implemented in ISL, and to expose a simple 1D tiling (with a single integer as its argument) we convert the tile size into the a single valued ISL isl_multi_val object.
Loop interchange is can easily be implemented by: (1) saving the partial schedule of the current node, (2) deleting the current node, and (3) inserting a new node with the saved schedule below the remaining node.
Fusing of two loops in a sequence (FUSE) is achieved by collecting the two filters (the children of the original sequence node) in an isl_union_set_list, then replacing the filters
of the two nodes by a single (union) filter, and finally inserting a new sequence node on top of the original sequence node.
To merge all band node under a sequence (FULL_FUSE), we take their partial schedules, intersect them with the corresponding filters and take the union. Then we introduce a new band node on top of the sequence101010Using isl_schedule_insert_partial_schedule!. and delete the original band nodes.
The 6 different SHIFT transformations, while simple on the surface, involve a tedious process involving the unpacking of the partial schedule, iterating through the sets of its domain and (re)constructing the adequate mupa111111Short for isl_multi_union_pw_aff!. object.
The SET_PARALLEL transformation simply insert a “mark” node, with an isl_id name "parallel" and the code generatation handles the rest.
Similarly to SET_PARALLEL, SET_LOOP_OPT is not strictly speaking a (polyhedral) transformation. It is implemented by simply calling the corresponding ISL function, and influences how code ISL generates code.
4.4.2. Extending Tadashi With New Transformations
The entry point of a transformation is a function with a SCoP index121212By convention size_t scop_idx!. as its first argument (potentially followed by other arguments) in ctadashi.cc. Unless the transformation is trivial, the main implementation is added to transformations.c as a function which takes a schedule node (with the rest of the arguments) and returns the transformed schedule node. This function is then called from the entry-point function with the current node of the SCoP specified by the SCoP index. The call needs to be prefixed a pre- and followed by a post-transform (convenience) functions, which take care of some bookkeeping to be able to rollback properly. On the Python side, a class derived from TranfromInfo stores the name of the entry-point and implements methods which check and provide information for the availability of the transformation on a given node. Finally the TrEnum and global TRANSFORMATION dictionary connect the TransformInfo into the mechanism implemented by the Node’s transform method.
4.5. Legality Checking
Flow of legality check for two examples.
After a transformation concludes, the legality check is invoked on the same node. The legality check implements the principles described in Section 2.2. Each step of the legality check in Algorithm 1 can be implemented with a single ISL function. Algorithm 1 is applied after every transformation with the exception of SET_PARALLEL. For the legality of parallel loops the entry in corresponding to the loop should be , and this can be achieved with a slight modification to Algorithm 1: we use the partial schedule instead of the complete schedule in steps 1 and 2, and in step 4 we compute and and check if they are both empty in step 5. Figure 4 shows two examples of the process of checking legality.
isl_union_map) from scop_info_t
scop_info_t
4.6. Code Generation
Code generation is based on the pet_loopback.c file from the PET repository, which extracts the polyhedral representation from an input C file, and generates the “same” code using the original schedule.
Apart from the obvious tweaking to use the modified schedule instead of the original one, and to make it cooperate with the bookkeeping setup we use in Tadashi’s back-end, the SET_PARALLEL transformation required additional modifications to the core code generation. Parallel loops are implemented as OpenMP parallel for loops, so the relevant pragma needs to be inserted. This is achieved by properly setting up the isl_ast_build and isl_ast_print_options objects, i.e., specifying custom after_each_mark and print_for callbacks for the AST build and print options object respectively.
5. Evaluation
We evaluate Tadashi using Random-RL, a mock ML model, verify the implementation and measure performace. All the measurements are done on a regular desktop with a 12 core AMD Ryzen 5 5600X 6-Core Processor and 64GiB of memory, running Arch Linux, with python 3.12.5-1, libisl 0.27-1, llvm 18.1.8-4 system packages, and Tadashi’s CMake script using the current master branch131313Git hash 824ab35b60bab089fb699e521e2e3bc6c07ce95a!. of PET141414From https://repo.or.cz/pet.git. tagged as pet 0.11.8.
5.1. Random-RL
To demonstrate the ease of use and performance of Tadashi, we show an example case of how Tadashi can be used in ML in an end-to-end fashion. We implement a mock RL script, Random-RL, described in Listings 3, with the agent randomly picking actions, which we evaluate on the Polybench (Reisinger, 2024) benchmark suit. Choosing a random action, consists of selecting a node in the schedule tree, a transformation, and parameters for that transformation. Technically, a random SCoP should also be chosen, but for simplicity, we assume that there is only one SCoP, which is conveniently true for the Polybench apps. The Random-RL agent keeps track of the last chosen node in the schedule tree, and the new node is selected by choosing randomly from its parent or one of it’s children (initially starting from the root). This is followed by a random sampling of all implemented transformations. The transformation might not be available for the selected node, so this process of picking random transformations and schedule nodes is repeated until a valid transformation-node pair is found. Finally, a valid set of transformation arguments is randomly picked, as described in Section 3.1.1.
5.2. Verification
With the Random-RL agent successfully sampling the transformations, we verify the correct implementation of Tadashi. First, all Polybench benchmarks are compiled and executed with the small dataset151515Using -DSMALL_DATASET!. and with the output saved.161616Using -DPOLYBENCH_DUMP_ARRAYS!. In an entirely automated manner, we generate 3 random transformations using the Random-RL’s agent (as described above), rollback the illegal ones, compile, run, and compare the output of the transformed program to the original. Post-inspect confirmed the outputs to be identical.
5.3. Performance
Breakdown of the results
Throughput
ML training algorithms need a vast number of code transformations, including the legality check and code generation, it would be informative to see how fast Tadashi performs these functions. With this in mind, we devise two experiments to measure a typical training cycle: one which involves a single polyhedral transformation and another with a sequence of 10 transformations (and legality checks), with a code generation at the end. A typical training cycle would cover functions performed by Tadashi, such as, extracting the polyhedral representation of the SCoPs, performing the transformation, checking the legality, generating code, as well as, compiling and running the program; steps with execution times which neither the user nor Tadashi can influence or circumvent. For both experiments, we use Polybench with the default dataset size, based on the reasoning that a smaller dataset would fit into cache memory and shrink the runtime to an extent where optimization would not be required.
The breakdown of the results can be found in Figure 5. The dominance of the runtime of the programs is obvious in more then half of the apps. As for the other benchmarks, with miniscule runtimes, we found that the functions performed by Tadashi always occupied less then half of the breakdown. Based on these numbers, using Tadashi should not be considerably more prohibitive than any other workflow which involves compiling and/or running the program.
To provide a better insight into Tadashi’s performance, we also calculate throughput from the breakdown data, which is shown in Figure 6. Only the number of transformations and legality checks are counted, since it can happen more often than SCoP extraction and code generation, and the latter two would also be accompanied by compilation and program execution, which, as we saw in Figure 5, would dominate the runtime. The results show Tadashi capability of sustaining 100s of iterations per second, which would be very practical for production ML solutions where 100,000s 100,000,000s of training examples or direct evaluations in RL would be the norm.
6. Conclusion
In this work we aim to lower the barrier for using ML in code auto generation, with guarantees on correctness. We rely on the polyhedral methods to generate code based on compositional loop transformations with mathematically proven guarantee of functional correctness. To this end we introduce Tadashi, a library which exposes an API simplified to the polyhedral methods, targeting the ML engineers and researchers not familiar with it. By using Tadashi, ML practitioners can design ML-driven methods based on the capability of Tadashi in sampling of the space of legal loop transformations.
Acknowledgements.
We are eternally grateful to Sven (skimo) Verdoolaege for his invaluable help by replying to our (often silly) questions on the ISL mailing list.References
- (1)
- Aho et al. (2007) Alfred V. Aho, Monica S. Lam, Ravi Sethi, and Jeffrey D. Ullman. 2007. Compilers: Principles, Techniques, & Tools (second edition ed.). Pearson/Addison Wesley, Boston. http://catdir.loc.gov/catdir/toc/ecip0618/2006024333.html
- Allamanis et al. (2024) Miltiadis Allamanis, Sheena Panthaplackel, and Pengcheng Yin. 2024. Unsupervised Evaluation of Code LLMs with Round-Trip Correctness. In ICML. OpenReview.net.
- Ashouri et al. (2018) Amir H. Ashouri, William Killian, John Cavazos, Gianluca Palermo, and Cristina Silvano. 2018. A Survey on Compiler Autotuning using Machine Learning. ACM Comput. Surv. 51, 5, Article 96 (sep 2018), 42 pages. https://doi.org/10.1145/3197978
- Baghdadi et al. (2021) Riyadh Baghdadi, Massinissa Merouani, Mohamed-Hicham Leghettas, Kamel Abdous, Taha Arbaoui, Karima Benatchba, and Saman P. Amarasinghe. 2021. A Deep Learning Based Cost Model for Automatic Code Optimization. In Proceedings of the Fourth Conference on Machine Learning and Systems, MLSys 2021, Virtual, April 5-9, 2021, Alex Smola, Alex Dimakis, and Ion Stoica (Eds.). mlsys.org. https://proceedings.mlsys.org/paper_files/paper/2021/hash/d9387b6d643efb25132be36f7b908d96-Abstract.html
- Beer et al. (2024) Robin Beer, Alexander Feix, Tim Guttzeit, Tamara Muras, Vincent Müller, Maurice Rauscher, Florian Schäffler, and Welf Löwe. 2024. Examination of Code generated by Large Language Models. arXiv:2408.16601 [cs.SE] https://arxiv.org/abs/2408.16601
- Ben-Nun et al. (2018) Tal Ben-Nun, Alice Shoshana Jakobovits, and Torsten Hoefler. 2018. Neural Code Comprehension: A Learnable Representation of Code Semantics. In Proceedings of the 32nd International Conference on Neural Information Processing Systems (NIPS’18). Curran Associates Inc., Red Hook, NY, USA, 3589–3601.
- Bondhugula et al. (2008) Uday Bondhugula, Albert Hartono, J Ramanujam, and P Sadayappan. 2008. A Practical Automatic Polyhedral Parallelizer and Locality Optimizer. (2008), 13.
- Cai et al. (2023) Zonghui Cai, Xiao Yang, MengChu Zhou, Zhi-Hui Zhan, and Shangce Gao. 2023. Toward explicit control between exploration and exploitation in evolutionary algorithms: A case study of differential evolution. Information Sciences 649 (2023), 119656. https://doi.org/10.1016/j.ins.2023.119656
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, Alex Ray, Raul Puri, Gretchen Krueger, Michael Petrov, Heidy Khlaaf, Girish Sastry, Pamela Mishkin, Brooke Chan, Scott Gray, Nick Ryder, Mikhail Pavlov, Alethea Power, Lukasz Kaiser, Mohammad Bavarian, Clemens Winter, Philippe Tillet, Felipe Petroski Such, Dave Cummings, Matthias Plappert, Fotios Chantzis, Elizabeth Barnes, Ariel Herbert-Voss, William Hebgen Guss, Alex Nichol, Alex Paino, Nikolas Tezak, Jie Tang, Igor Babuschkin, Suchir Balaji, Shantanu Jain, William Saunders, Christopher Hesse, Andrew N. Carr, Jan Leike, Josh Achiam, Vedant Misra, Evan Morikawa, Alec Radford, Matthew Knight, Miles Brundage, Mira Murati, Katie Mayer, Peter Welinder, Bob McGrew, Dario Amodei, Sam McCandlish, Ilya Sutskever, and Wojciech Zaremba. 2021. Evaluating Large Language Models Trained on Code. arXiv (2021). arXiv:2107.03374 [cs.LG]
- Coello (2022) Carlos A. Coello Coello. 2022. Constraint-handling techniques used with evolutionary algorithms. In Proceedings of the Genetic and Evolutionary Computation Conference Companion (Boston, Massachusetts) (GECCO ’22). Association for Computing Machinery, New York, NY, USA, 1310–1333. https://doi.org/10.1145/3520304.3533640
- Cotroneo et al. (2024) Domenico Cotroneo, Alessio Foggia, Cristina Improta, Pietro Liguori, and Roberto Natella. 2024. Automating the correctness assessment of AI-generated code for security contexts. Journal of Systems and Software 216 (2024), 112113. https://doi.org/10.1016/j.jss.2024.112113
- Cummins et al. (2020) Chris Cummins, Zacharias V. Fisches, Tal Ben-Nun, Torsten Hoefler, and Hugh Leather. 2020. ProGraML: Graph-based Deep Learning for Program Optimization and Analysis. (March 2020). https://doi.org/10.48550/arXiv.2003.10536
- Cummins et al. (2023) Chris Cummins, Volker Seeker, Dejan Grubisic, Mostafa Elhoushi, Youwei Liang, Baptiste Roziere, Jonas Gehring, Fabian Gloeckle, Kim Hazelwood, Gabriel Synnaeve, and Hugh Leather. 2023. Large Language Models for Compiler Optimization. arXiv:2309.07062 [cs.PL]
- Cummins et al. (2024) Chris Cummins, Volker Seeker, Dejan Grubisic, Baptiste Roziere, Jonas Gehring, Gabriel Synnaeve, and Hugh Leather. 2024. Meta Large Language Model Compiler: Foundation Models of Compiler Optimization. https://doi.org/10.48550/arXiv.2407.02524 arXiv:2407.02524 [cs]
- Dehaerne et al. (2022a) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022a. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. https://doi.org/10.1109/ACCESS.2022.3196347
- Dehaerne et al. (2022b) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022b. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. https://doi.org/10.1109/ACCESS.2022.3196347
- Du et al. (2024) Xueying Du, Mingwei Liu, Kaixin Wang, Hanlin Wang, Junwei Liu, Yixuan Chen, Jiayi Feng, Chaofeng Sha, Xin Peng, and Yiling Lou. 2024. Evaluating Large Language Models in Class-Level Code Generation. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 81, 13 pages. https://doi.org/10.1145/3597503.3639219
- Feautrier (1991) Paul Feautrier. 1991. Dataflow Analysis of Array and Scalar References. (1991), 37. Introduction of SCoP.
- Feautrier (1996) Paul Feautrier. 1996. Some Efficient Solutions to the Affine Scheduling Problem Part I One-dimensional Time. International Journal of Parallel Programming 21 (Aug. 1996). https://doi.org/10.1007/BF01407835
- Feautrier (1997) Paul Feautrier. 1997. Some Efficient Solutions to the Affine Scheduling Problem Part II Multidimensional Time. International Journal of Parallel Programming 21 (Jan. 1997). https://doi.org/10.1007/BF01379404
- Github ([n. d.]) Github. [n. d.]. GitHub Copilot Your AI Pair Programmer. https://github.com/features/copilot
- Grosser et al. (2012) Tobias Grosser, Armin Groesslinger, and Christian Lengauer. 2012. Polly - Performing Polyhedral Optimizations on a Low-Level Intermediate Representation. Parallel Process. Lett. 22, 04 (Dec. 2012), 1250010. https://doi.org/10.1142/S0129626412500107
- Ivanov et al. (2024) Ivan R. Ivanov, Joachim Meyer, Aiden Grossman, William S. Moses, and Johannes Doerfert. 2024. Input-Gen: Guided Generation of Stateful Inputs for Testing, Tuning, and Training. arXiv:2406.08843 [cs.SE] https://arxiv.org/abs/2406.08843
- Jiang et al. (2024) Juyong Jiang, Fan Wang, Jiasi Shen, Sungju Kim, and Sunghun Kim. 2024. A Survey on Large Language Models for Code Generation. arXiv:2406.00515 [cs.CL]
- Kind et al. (2022) Cecilia Conde Kind, Michael Canesche, and Fernando Magno Quintao Pereira. 2022. Jotai: a Methodology for the Generation of Executable C Benchmarks. Technical Report. Technical Report 02-2022. Universidade Federal de Minas Gerais.
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Proceedings of the 34th International Conference on Neural Information Processing Systems (NIPS ’20). Curran Associates Inc., Red Hook, NY, USA, 9459–9474.
- Liang et al. (2023) Youwei Liang, Kevin Stone, Ali Shameli, Chris Cummins, Mostafa Elhoushi, Jiadong Guo, Benoit Steiner, Xiaomeng Yang, Pengtao Xie, Hugh Leather, and Yuandong Tian. 2023. Learning compiler pass orders using coreset and normalized value prediction. In Proceedings of the 40th International Conference on Machine Learning (Honolulu, Hawaii, USA) (ICML’23). JMLR.org, Article 855, 17 pages.
- Mankowitz et al. (2023) Daniel J. Mankowitz, Andrea Michi, Anton Zhernov, Marco Gelmi, Marco Selvi, Cosmin Paduraru, Edouard Leurent, Shariq Iqbal, Jean-Baptiste Lespiau, Alex Ahern, Thomas Köppe, Kevin Millikin, Stephen Gaffney, Sophie Elster, Jackson Broshear, Chris Gamble, Kieran Milan, Robert Tung, Minjae Hwang, Taylan Cemgil, Mohammadamin Barekatain, Yujia Li, Amol Mandhane, Thomas Hubert, Julian Schrittwieser, Demis Hassabis, Pushmeet Kohli, Martin Riedmiller, Oriol Vinyals, and David Silver. 2023. Faster Sorting Algorithms Discovered Using Deep Reinforcement Learning. Nature 618, 7964 (June 2023), 257–263. https://doi.org/10.1038/s41586-023-06004-9
- Moses et al. (2021) William S Moses, Ruizhe Zhao, Lorenzo Chelini, and Oleksandr Zinenko. 2021. Polygeist: Affine C in MLIR. (2021), 12.
- Munoz et al. (2024) Michael Munoz, Juan Torre, Stefany Lopez, Stiven Herrera, and Christian Uribe. 2024. Comparative Study of AI Code Generation Tools: Quality Assessment and Performance Analysis. LatIA 2 (07 2024), 104. https://doi.org/10.62486/latia2024104
- Mustafa (2022) Dheya Mustafa. 2022. A Survey of Performance Tuning Techniques and Tools for Parallel Applications. IEEE Access 10 (2022), 15036–15055. https://doi.org/10.1109/ACCESS.2022.3147846
- Phan (2019) Hung Phan. 2019. Self Learning from Large Scale Code Corpus to Infer Structure of Method Invocations. https://arxiv.org/abs/1909.03147v1
- Pouchet et al. (2011) Louis-Noël Pouchet, Uday Bondhugula, Cédric Bastoul, Albert Cohen, J. Ramanujam, P. Sadayappan, and Nicolas Vasilache. 2011. Loop transformations: convexity, pruning and optimization. In Proceedings of the 38th Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages (Austin, Texas, USA) (POPL ’11). Association for Computing Machinery, New York, NY, USA, 549–562. https://doi.org/10.1145/1926385.1926449
- Reisinger (2024) Matthias Reisinger. 2024. MatthiasJReisinger/PolyBenchC-4.2.1. https://github.com/MatthiasJReisinger/PolyBenchC-4.2.1
- Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessi, Roberta Raileanu, Maria Lomeli, Eric Hambro, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language Models Can Teach Themselves to Use Tools. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 68539–68551. https://proceedings.neurips.cc/paper_files/paper/2023/file/d842425e4bf79ba039352da0f658a906-Paper-Conference.pdf
- Shido et al. (2019) Yusuke Shido, Yasuaki Kobayashi, Akihiro Yamamoto, Atsushi Miyamoto, and Tadayuki Matsumura. 2019. Automatic Source Code Summarization with Extended Tree-LSTM. In 2019 International Joint Conference on Neural Networks (IJCNN). 1–8. https://doi.org/10.1109/IJCNN.2019.8851751
- Silver et al. (2016) David Silver, Aja Huang, Chris J. Maddison, Arthur Guez, Laurent Sifre, George van den Driessche, Julian Schrittwieser, Ioannis Antonoglou, Veda Panneershelvam, Marc Lanctot, Sander Dieleman, Dominik Grewe, John Nham, Nal Kalchbrenner, Ilya Sutskever, Timothy Lillicrap, Madeleine Leach, Koray Kavukcuoglu, Thore Graepel, and Demis Hassabis. 2016. Mastering the game of Go with deep neural networks and tree search. Nature 529, 7587 (01 Jan 2016), 484–489. https://doi.org/10.1038/nature16961
- Stein et al. (2020) Aviel J. Stein, Levi Kapllani, Spiros Mancoridis, and Rachel Greenstadt. 2020. Exploring Paraphrasing Techniques on Formal Language for Generating Semantics Preserving Source Code Transformations. In 2020 IEEE 14th International Conference on Semantic Computing (ICSC). 242–248. https://doi.org/10.1109/ICSC.2020.00051
- Sutton and Barto (2018) Richard S Sutton and Andrew G Barto. 2018. Reinforcement Learning: An Introduction. MIT press.
- Trümper et al. (2023) Lukas Trümper, Tal Ben-Nun, Philipp Schaad, Alexandru Calotoiu, and Torsten Hoefler. 2023. Performance Embeddings: A Similarity-Based Transfer Tuning Approach to Performance Optimization. In Proceedings of the 37th ACM International Conference on Supercomputing (Orlando, FL, USA) (ICS ’23). Association for Computing Machinery, New York, NY, USA, 50–62. https://doi.org/10.1145/3577193.3593714
- Vasilache et al. (2018) Nicolas Vasilache, Oleksandr Zinenko, Theodoros Theodoridis, Priya Goyal, Zachary DeVito, William S. Moses, Sven Verdoolaege, Andrew Adams, and Albert Cohen. 2018. Tensor Comprehensions: Framework-Agnostic High-Performance Machine Learning Abstractions. CoRR abs/1802.04730 (2018). arXiv:1802.04730 http://arxiv.org/abs/1802.04730
- Verdoolaege (2010) Sven Verdoolaege. 2010. Isl: An Integer Set Library for the Polyhedral Model. In Mathematical Software – ICMS 2010, Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama (Eds.). Springer, Berlin, Heidelberg, 299–302. https://doi.org/10.1007/978-3-642-15582-6_49
- Verdoolaege et al. (2013) Sven Verdoolaege, Juan Carlos Juega, Albert Cohen, José Ignacio Gómez, Christian Tenllado, and Francky Catthoor. 2013. Polyhedral Parallel Code Generation for CUDA. ACM Trans. Archit. Code Optim. 9, 4 (Jan. 2013), 54:1–54:23. https://doi.org/10.1145/2400682.2400713
- Verdoolaege and Grosser (2012) Sven Verdoolaege and Tobias Grosser. 2012. Polyhedral Extraction Tool. (2012), 8.
- Verdoolaege et al. (2014) Sven Verdoolaege, Serge Guelton, Tobias Grosser, and Albert Cohen. 2014. Schedule Trees. In International Workshop on Polyhedral Compilation Techniques. Vienna, Austria. https://lirias.kuleuven.be/1656334
- Verdoolaege et al. ([n. d.]) Sven Verdoolaege, Polly Labs, and KU Leuven. [n. d.]. Presburger Formulas and Polyhedral Compilation. ([n. d.]), 174.
- Wang et al. (2024) Ruotong Wang, Ruijia Cheng, Denae Ford, and Thomas Zimmermann. 2024. Investigating and Designing for Trust in AI-powered Code Generation Tools. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (Rio de Janeiro, Brazil) (FAccT ’24). Association for Computing Machinery, New York, NY, USA, 1475–1493. https://doi.org/10.1145/3630106.3658984
- Wei et al. (2024) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2024. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
- While and Hingston (2013) Lyndon While and Philip Hingston. 2013. Usefulness of infeasible solutions in evolutionary search: An empirical and mathematical study. In 2013 IEEE Congress on Evolutionary Computation. 1363–1370. https://doi.org/10.1109/CEC.2013.6557723
- Zhou et al. (2023) Shuyan Zhou, Uri Alon, Frank F. Xu, Zhengbao Jiang, and Graham Neubig. 2023. DocPrompting: Generating Code by Retrieving the Docs. In ICLR. OpenReview.net.