### v2

# Tadashi: Enabling AI-Based Automated Code Generation With Guaranteed Correctness

Abstract.
Frameworks and domain-specific languages for auto-generating code have traditionally depended on human experts to implement rigorous methods ensuring the legality of code transformations. Recently, machine learning (ML) has gained traction for generating code optimized for specific hardware targets. However, ML approaches-particularly black-box neural networks-offer no guarantees on the correctness or legality of the transformations they produce. To address this gap, we introduce Tadashi, an end-to-end system that leverages the polyhedral model to support researchers in curating datasets critical for ML-based code generation. Tadashi provides an end-to-end system capable of applying, verifying, and evaluating candidate transformations on polyhedral schedules with both reliability and practicality. We formally prove that Tadashi guarantees the legality of generated transformations, demonstrate its low runtime overhead, and showcase its broad applicability. Tadashi available at https://github.com/vatai/tadashi/.
Tadashi sampling the space of transformations
1. Introduction
Machine learning (ML) is disrupting all forms of digital media, and the creation of computer programs is no exception. ML can be and has been applied in various ways with the aim of making the work of programmers (Dehaerne et al., 2022) easier: generating code from “natural languages” (Phan, 2019), write a summary (e.g., write commit messages) of a piece of code (Shido et al., 2019), auto-complete code on-the-fly while it’s being written (Github, 2022) among others. In this paper, we consider the task of optimising code to a hardware target, i.e., using ML to transform the hotspots in the code, i.e., loops, in a way that improves performance.
Checking if the program generated by the ML model is faster then the baseline is just a matter of running it, then comparing against the run-time of the original code. However, it is a major challenge to get formal guarantees that the transformed code produces the same output as the original (Dehaerne et al., 2022; Munoz et al., 2024; Jiang et al., 2024). This impedes the adoption of ML for code generation and transformation in mission critical scientific/engineering codes, or codes that come with high liability, such as in the financial, legal, and corporate sectors. Additionally, there is a lack of tooling for generating and sharing parameterized datasets comprised of legal auto-generated code, which further constraints progress in the field since, to date, ML-based solutions have to curate their datasets.
To this end, we developed Tadashi, a Python library for sampling the search space of possible loop transformations, with the capability to verify the legality (i.e., correctness) of the sampled transformations. Tadashi allows the user to programmatically specify the pattern for sampling legal individual or combinations of transformations, based on the preference of the ML method that would be used to explore the search space of transformations. Through the use of the polyhedral model, Tadashi is capable of enabling ML engineers to explore novel and more aggressive methods of code optimisation in the datasets they curate, while guaranteeing that the results of the generated code are functionally identical to the original program. Tadashi is an end-to-end solution to transform, verify and evaluate transformations, emphasizing the ease-of-use for ML practitioners, with support for HPC oriented libraries such as MPI4Py futures (Rogowski et al., 2023); Tadashi aims to provide a simple Python interface, curating datasets or performing direct code-generation at scale to be used with different ML approaches.
We envision Tadashi with the overarching purpose of helping clusters and supercomputers reach their full potential. Namely, after the deployment of a new (or existing) supercomputing system, one can train an ML model to learn to transform hotspots in scientific software to run optimally on the given hardware. By basing their ML models on Tadashi, end users will have both fully correct and highly optimized code. With this in mind Tadashi would be a viable and practicable option when porting existing science codes to new supercomputers. This is particularly a huge challenge, as demonstrate in recent history by the Exascale Computing Project (Heroux, 2024; exa, [n. d.]), which required 7-years, $1.8 billion, and approximately 1,000 researchers from DOE national laboratories to port the codes to new systems (among other goals of ECP).
In this paper, we make the following contributions:
We introduce Tadashi, a Python-based end-to-end system designed for ML practitioners seeking to design tools for code auto generation with correctness guarantees. Tadashi lowers the barrier to entry of polyhedral loop transformations; a model that can give mathematically provable guarantees on the correctness (i.e., legality) of the transformed code.
We demonstrate how Tadashi can be used by an ML system by implementing two representative ML models on top of Tadashi.
We show the capability of Tadashi to speed up software on Polybench suite and two science applications.
2. Background and Motivation
In this section we discuss the motivation for designing a library that leverages the polyhedral model as bases for guaranteeing correctness of ML-driven code generation.
2.1. Correctness in ML-Based Code Auto-Generation
ML is being increasingly used in code auto-generation, at different levels of abstraction, e.g., source code (Stein et al., 2020), abstract syntax tree (AST) (Shido et al., 2019), polyhedral (Baghdadi et al., 2021), dependency graphs (Cummins et al., 2020), intermediate Representations (IR) (Ben-Nun et al., 2018), and assembly instructions (Mankowitz et al., 2023).
To ensure correctness, existing approaches for using ML in code generation or transformation rely on one of the following approaches:
Extensive unit testing: for this, widely used, approach, a suite of unit tests is used to validate correctness (Cummins et al., 2024; Beer et al., 2024; Du et al., 2024; Chen et al., 2021). While unit testing can be effective, it provides no guarantees on correctness111”Program testing can be used to show the presence of bugs, but never to show their absence!”– Edsger W. Dijkstra.
In summary, the approaches above either provide no formal guarantee on functional correctness, or confine the user to a very constrained form of code generation. While some have suggested there is a tradeoff between correctness and performance, (Ouyang et al., 2024) we take a different view—emphasizing that correctness remains paramount. This perspective is reinforced by recent findings (202, 2025) where what was initially reported as a 100x speedup from an AI-generated solution was, upon closer examination, actually a 3x slowdown. In addition, we observed that for the above approaches, it was the burden of the researchers to implement a sampling strategy and curate the dataset. Solving those two problems, correctness and agency, is the motivation for Tadashi.
2.2. The Polyhedral Model
Polyhedral techniques have a long history with one of the seminal papers dating back to 1996 (Feautrier, 1996) and are well established and used in compilers (Aho et al., 2007). The polyhedral model is a very efficient way to manipulate and reason about nested loops, including the capability to check the legality222Legality is the technical term, used in the polyhedral literature with a precise definition which translates to the transformed code calculating the same results as the original, i.e., functional correctness. of the transformations. It has been used in optimization pass of compilers (Grosser et al., 2012), as well as standalone source-to-source optimization tools (Bondhugula et al., 2008; Verdoolaege et al., 2013) relying on integer linear programming with some optimization heuristic.
The polyhedral model is a representation for loop nests with regular boundaries and memory accesses, know as static control parts (see Listing 1 an example and Section 2.3 for a description). The polyhedral representation separates the statement instances from the order in which they are executed: Each of these instances is then mapped to an integer tuple, a timestamp, and then the statement instances are ordered according the lexicographical order of their timestamps. For a given statement , the set of statement instances is called the domain and is denoted as ; while the mapping which determines their order is the schedule, which is denoted by . Listing 1 and Figure 2(d) give a simple example of a SCoP and possible schedules.
The main idea of polyhedral optimization is to transform the schedule (or create a new one from scratch) which reorders the statement instances so they are executed more efficiently on a given hardware (e.g., by increasing memory locality). As one could expect, not all schedule transformations are legal, however checking legality is very convenient in the polyhedral model. Dependency analysis (Feautrier, 1991) yields a directed dependecy graph . The vertices of are statements (or equivalently their domains), and the edges are labelled with maps which describe the dependencies on the level of statement instances such as
Applying the schedule to the dependee and dependent of these maps takes them into the time domain, and (1) becomes ( and are timestamps). If lexicographically precedes , i.e.,
then the schedule is legal, because it maps the dependent to a later time then the timestep of the dependee. Because (2) is equivalent to , we can (symbolically) compute and if is lexicographically positive, the schedule is legal. Computing is useful because it provides us with even more useful information: if an entry in is zero, it means there are no dependencies along the corresponding dimension of the schedule, and consequently the generated loop for that dimension can be parallelized.
Example schedules
2.3. Constraints of the Polyhedral Model
To be able to represent the domain, schedule, and dependencies efficiently, the source codes represented by the polyhedral model are restricted to Static Control Parts, or SCoPs for short (Listing 1 is a simple example). While these restrictions are not negligible, most of high performance and science codes do not violate them. The main restrictions of SCoPs are the following: (1) The control-flow must be static, which means it does not depend on the input, other than parameters (such as loop boundaries) which remain constant during execution, (2) there is no memory aliasing, no pointer manipulation, and (3) the symbolic expressions and formulas, used throughout the model, need to be expressed using a matrix and a vector as a (quasi-)affine expression, i.e., (sometimes referred to as Presburger formulas (Verdoolaege et al., 2021)). For example, the domains are expressed as ; schedules are (quasi-)affine functions . The polyhedral model can also handle tiling, so the formulas include integer division with a constant as well, and this is what the “quasi-” prefix refers to. However, for brevity and clarity, we omit discussion of these details and for a more comprehensive description of the polyhedral model we recommend (Feautrier, 1996, 1997; Verdoolaege et al., 2021).
Abiding by those restrictions, which is reasonably possible in a wide range of high performance science and engineering applications, enables ML systems that use Tadashi to guarantee correct and legal output at a scale, unlike most other existing solutions.
3. Tadashi: A Library to Sample Legal Polyhedral Transformations
Tadashi is a library that provide the means for ML practitioners to sample the space of legal polyhedral loop transformations. The user can use Tadashi to describe a sampling strategy that is suited for the ML approach, e.g., uniform sampling to curate datasets vs. guided exploration of the space of legal transformations. Figure 1 gives an overview of Tadashi. At an architectural level, Tadashi consists of a Python front-end and a C/C++ back-end (aptly named ctadashi). The front-end provides a minimal and easy-to-use interface to sample from the space of possible transformations, check their legality, and generate the transformed code. More importantly, the front-end being in Python encourages fast adoption by ML practitioners since Python is the de facto language used in AI/ML frameworks. The back-end implements this functionality, using the Integer Set Library (ISL) (Verdoolaege, 2010) and Polyhedral Extract Tool (PET) (Verdoolaege and Grosser, 2012) polyhedral libraries.
Existing ISL wrapper libraries for Python expose the full API of the library, which can be daunting for users not versed in the concepts of polyhedral compilation. The goal of Tadashi is to lower this barrier to entry for ML engineers and researchers by providing out-of-the-box implementations of transformations, legality check and code generation. Moreover, the functionality is exposed in a more idiomatically Pythonic way. For example, a tree of nested loops is presented is a hierarchy of nested objects similar to how HTML parse trees are used in Python libraries, i.e., a transformation on a given node can be done by calling an appropriate method of a given object, while hiding all bookkeeping necessary for interacting with ISL and PET libraries.
3.1. Python Front-End: Polyhedral Made Easy
To demonstrate the ease of use of Tadashi, Listing 2 performs a loop fission, checks legality, generates new code and measures speedup in no more than 10 lines of a fully functional Python example. Line 3 creates the app object, which stores the information required by Tadashi to load the input file. The location of the transformation is determined by selecting a node objects on line 4. The fission transformation is performed on line 5. The new code code is generated in line 6.
To pass an input file to Tadashi, it must be wrapped in an instance of the App abstract base class. Currently Tadashi provides two implementations of App: the Simple class which can be used for simple single-file programs and Polybench (used in line 3) to load Polybench kernels (Reisinger, 2024). The classes derived from App store the input filename to Tadashi and other information needed by Tadashi when parsing the file. For convenience, App also defines compile() and measure() methods.
The polyhedral representation of all loop nests (SCoPs) are implemented as an array-like scops object (line 4), and from the first SCoP (scops[0]) a node of the schedule trees (Verdoolaege et al., 2014) is selected. For a user not familiar with the polyhedral model and schedule trees, a node is “an entity that can be transformed”, which usually means a loop in the loop nest (a brief discussion of schedule trees can be found in Section 4). Since some transformations are not available for certain loops (e.g., it doesn’t make sense to apply 3D tiling on a 2-deep loop nest), the Node class provides methods to query the available transformations and their (valid) arguments. The nodes of a scop can be accessed through the (dynamic) scop.schedule_tree[] as a list, or as a tree using the goto_child(child_idx) and goto_parent() methods of nodes.333The root is scop.schedule_tree[0]!.
A code transformation is performed by passing a TrEnum (an Enum type enumerating all transformations) object to the node’s transform() method (with the transformation-specific argument). A summary of implemented transformations (and their parameters) can be found in Table 1. TILE1D, (TILE2D and TILE3D) implements 1D, (2D and 3D) tiling (higher dimentional tiling can be expressed tiling and INTERCHANGE transfromations). INTERCHANGE switches an outer loop with the loop directly under it. For PARTIAL shift transformations, the statement index (stmt idx) selects only one statement to be shifted, while the FULL versions shift all statements. The VAL shift transformation shift the schedule by a constant, the VAR and PARAM versions shift by an iterator (iter idx) of an outer loop or a parameter (param idx), multiplied by a coefficient (coeff). FULL_FUSE fuses all loops in a sequence, while FUSE fuses two selected loops. Similarly, FULL_SPLIT and SPLIT perform loop fission. SET_PARALLEL and SET_LOOP_OPT modify how code of a loop is generated. These are primitive transformations, in the sense that they are composable, so we can apply a sequence of transformations on the same SCoP. In the current version of Tadashi, we prioritized the the implementation of loop transformations that have high impact on performance (Pouchet et al., 2011).
The return value of the transform() method is the result of the legality check and, based on it, the users can decide what they want to do. In case of a legal transformation, one likely action would be generating the C code (line 6 in Listing 2), compiling it (line 8) and measuring the execution time (line 9), while for an illegal transformation one could invoke the rollback() method of the node, to return the polyhedral representation to the previous, legal state. Additionally, Scop objects provide a reset method and a method to apply a list of transformations, useful to implement “arbitrary rollback”.
TrEnum |
Args |
TILE |
tile size |
INTERCHANGE |
FUSE |
2 loop indices |
FULL_FUSE |
SPLIT |
1 stamt index |
FULL_SPLIT |
FULL_SHIFT_VAL |
const shift value |
PARTIAL_SHIFT_VAL |
stmt idx, const shift value |
FULL_SHIFT_VAR |
coeff, shift iter idx |
PARTIAL_SHIFT_VAR |
stmt idx, coeff, shift iter idx |
FULL_SHIFT_PARAM |
coeff, shift param idx |
PARTIAL_SHIFT_PARAM |
stmt idx, coeff, shift param idx |
SCALE |
coeff |
SET_PARALLEL |
SET_LOOP_OPT |
AstLoopType enum |
3.1.1. Sampling Strategies
Listing 2 demonstrates the application of a single primitive transformation. However, in consequence of the compositionality of polyhedral transformations, the user can apply a sequence of transformations and legality checks to a SCoP (before generating the code) and hence obtain a different composite transformation. This is a new sample from the space of polyhedral transformations, different from its component primitive transformations. Put differently, sampling the space of polyhedral transformations with Tadashi consists of applying a sequence of transformations to a SCoP, where each entry consists of specifying a location (a node in the schedule tree), the transformation (TrEnum) and a set of arguments. A location can be selected using the schedule_tree[] member or with a combination of goto_* methods of the Node class. The transformation can be selected from the possible values TrEnum values and checked if it is available on the selected node. Finally, selecting an argument is somewhat convoluted, since the available values depend not only on the selected transformation, but also on the concrete properties of the selected schedule node (such as its partial schedule, type and number of children etc.). The get_args() method of Nodes simplifies this.
3.2. Using Different ML Methods With Tadashi
The goal of Tadashi is to lower the barrier to entry for user-guided sampling of legal loop transformations. Tadashi does not in itself implement any type of ML; we do however design Tadashi with considerations to how various ML approaches can use it. The design of Tadashi supports the following ML search space exploration approaches:
3.2.1. Reinforcement Learning (RL)
RL (Sutton and Barto, 2018) is a prime candidate to use Tadashi, since Tadashi’s interface maps naturally to RL concepts. Similar to choosing winning moves in a game, an RL agent chooses promising transformations. The transformed code is executed by the runtime and speedup over the original version serves a the reward signal. Sequences of transformations can be sampled by using approaches such as Monte Carlo tree search (MCTS), where a neural network model gives scores to possible choices - an approach similar to Alpha Go(Silver et al., 2016) system.
3.2.2. Supervised Learning
Different versions of transformed source codes and corresponding speedups can be saved in a database. A model to predict speedups from transformations can be trained in supervised fashion to be used as a value model in an off-policy reinforcement learning setup or in other applications.
Such approach is especially valuable, because evaluation of transformed code takes significantly more time, compared to selecting and performing transformations (see Section 5.4).
The choice of transformations can be done in random or by grid search. Many transformations require additional corresponding parameters, e.g., choosing to tile a loop requires choosing a tile size. These arguments (second column in Table 1) allow the user to combinatorially increase the coverage of the search space to consequently increase the coverage of the curated dataset. For example, starting with a limited set of input C source files containing loop nests, one could run Tadashi to select random transformations (on a random input file, on random nodes of the schedule tree, with random parameters), then store the new code with the result of the legality check and the measured execution time. The dataset generated this way would consist of (input, output) pairs, where the input is (the name of) a C source file, and the output consists of: (1) the transformation, (2) transformed code, (3) legality of the transformation, (4) execution time of the transformed program.
In the case of legal transformations, the transformed code can be added to the pool of input files, expanding the number of different input files ad infinitum.
Another, approach is to sample sequences of transformations, i.e., generate a sequence (of a random length) of random transformations, where the individual transformations are sampled similarly as described in the previous paragraph. The legality check, code generation and performance measurement are performed only at the end of the sequence.
With basic Python programming, e.g., using the itertools package, one can easily generate a regular “grid” of transformations, which is adequate for scenarios similar to auto-tuning.
3.2.3. Evolutionary Computation Algorithms
Similar to RL, evolutionary computation (EC) algorithms use a mix of exploration and exploitation to traverse the search space, however without relying on learnable parameters and performing search using more direct heuristics (Cai et al., 2023). Exploration refers to searching for new, diverse solutions by exploring a wide range of the solution space, while exploitation focuses on refining and improving existing solutions by intensively searching around the best solutions found so far. As mentioned above in other ML approaches, Tadashi’s sampling of the search space can be parameterized by the user to explore a wide range of the search space (exploration) or intensive search in a narrow region of the search space (exploitation). The runtime of the transformed code variants would be used as the objective function.
The exploration capability of EC algorithms can in some instances improve if infeasible solutions appear in the population (While and Hingston, 2013), and then later be penalized to disappear over time from the population (Coello, 2022). Tadashi internally manipulates the schedule to generate both legal and illegal transformations, where we then filter out the illegal transformations. Since illegal transformations can be of value for the exploration capability of EC algorithms, we allow the user to sample illegal transformations simply by not rolling them back and generating code disregarding the result of the legality check.
3.2.4. With LLM Agents
The most performant transformations identified by any of the aforementioned methods can be used as training data for Large Language Models for better end-to-end code generation or as additional signal to improve overall capacity of the model (Du et al., 2024).
However, in line with the motivation for this paper, an even more promising application of Tadashi would be to be called directly by an LLM agent in a “tool calling” paradigm (Schick et al., 2023), similar to Retrieval Augmented Generation (RAG) systems (Lewis et al., 2020) or Chain of Thought approach (Wei et al., 2024).
3.2.5. Auto-Tuning
For practical reasons, ML approaches used in auto-tuning (Mustafa, 2022; Ashouri et al., 2018) mostly rely on local search methods with focused search around particular regions in the search space, after aggressively pruning the search space. Similar to the exploitation in EC algorithms, Tadashi can be setup to narrow down the search around the region for auto-tuning, based on the post-pruning range of interest.
3.2.6. ML-driven deployment of applications on new hardware
In addition to supporting the usage patterns of the ML approaches listed above, Tadashi can drive an ML approach to learnhardware specific optimisations on new hardware. After the training phase, applications ported to the new hardware can do inference on the ML model, leveraging Tadashi, to optimize the applications for the new hardware.
4. Implementation of Tadashi
We now revisit the end-to-end example of Listing 2, but with our attention directed to implementation details and what is happening under the hood, in the ctadashi back-end depicted in Figure 3.
Overview of the ctadashi back-end.
4.1. The App Class
The App class (Listing 2 line 3) is completely written in Python and provided mostly for convenience. It’s main functions are to provide the source_path and include_paths properties, the former being the C source file which will be transformed, and the later a list of paths used to set the required environment variable,444Namely C_INCLUDE_PATH!. which is necessary, because when PET runs a compiler pass on the input file, it will throw an error if it cannot find an included header file.
4.2. SCoPs
Accessing the Scops object (line 4) sets things in motion in the C/C++ back-end of Tadashi. The input file is passed through PET’s C source transformation function555Namely pet_transform_C_source. which detects all the SCoPs in the file, but instead of doing any transformations, the pet_scop objects are used to populate a container tadashi_scop structs. These structs store bookkeeping information and ISL objects, such as the domain, schedule and dependencies, required for the legality checks. Dependencies are computed by ISL as described in (Verdoolaege et al., 2021) and algorithm adopted from PPCG (Verdoolaege et al., 2013) is executed at this point to eliminate dead code. Tadashi also handles the management of loading multiple apps, requiring the creation, storage and deletion of multiple Scops.
4.3. Schedule Tree and Navigation
Transformations in Tadashi are performed on schedule trees, which are alternative ways to represent the domains and schedules in a SCoP (see Section 2.2). The leaves of the schedule tree are the statements, the root is the union of their domains. The loops surrounding a statement (in the source code) correspond to band nodes in the tree. On the path from the root to the leaf, each band node has a partial schedule and the complete schedule of the statement (i.e., leaf) can be constructed by concatenating the partial schedules. Other nodes in the schedule tree are sequence and filter nodes. Each sequence node has multiple filter nodes as children and together they represent a sequence of statements and/or nested loops. The sequence node is the placeholder node for the entire sequence, while its children act as roots, i.e., domains for the sub-schedule trees.
Each transformation in Tadashi is defined on a node of a schedule tree, so line 4 in Listing 2 involves the following mechanism. First, scops[i] returns a Python Scop object which stores the index of the corresponding tadashi_scop entry on the C++ side. This Scop object wraps all the C++ functions exposed to Python by calling them with a scop_idx parameter from an object method with same name. These functions include navigation operations666Namely goto_root(), goto_parent(), and goto_child(idx). to update the current node to point to its parent, its child, or to the root of the schedule tree (for the selected tadashi_scop object). The schedule_tree property on line 4 is dynamically generated, and exposes the schedule tree as a list of nodes by doing a depth-first traversal. During the traversal, we build a list of child indexes which enables Tadashi to locate any node by calling goto_child for each element in the list (implemented as the locate() method of a Node). In addition to the navigation methods, Tadashi exposes wrappers to query properties of nodes, (such as node type, number of children, partial schedule, parameters, enclosing iterators etc.), and so enables the Python Node object to emulate the ISL schedule tree and its nodes in C++.
4.4. Transformations
After updating the current node to point to the desired location, invoking node.transform() calls the C++ function implementing the corresponding transformation. The transformation is performed on the current node of the SCoP specified by the index obtained from the node (similar to the navigation and query functions from the previous paragraph). To be able to rollback the transformation, the current node is copied, and restored if the user calls the rollback method (usually after the transformation turns out to be illegal).
4.4.1. Implementation Details of Individual Transformations
Multi dimensional tiling is already implemented in ISL, and to expose a simple 1D tiling (with a single integer as its argument) we convert the tile size into the a single valued ISL isl_multi_val object.
Loop interchange is can easily be implemented by: (1) saving the partial schedule of the current node, (2) deleting the current node, and (3) inserting a new node with the saved schedule below the remaining node.
Fusing of two loops in a sequence (FUSE) is achieved by collecting the two filters (the children of the original sequence node) in an isl_union_set_list, then replacing the filters
of the two nodes by a single (union) filter, and finally inserting a new sequence node on top of the original sequence node.
To merge all band node under a sequence (FULL_FUSE), we take their partial schedules, intersect them with the corresponding filters and take the union. Then we introduce a new band node on top of the sequence777Using isl_schedule_insert_partial_schedule!. and delete the original band nodes.
The 6 different SHIFT transformations, while simple on the surface, involve a tedious process involving the unpacking of the partial schedule, iterating through the sets of its domain and (re)constructing the adequate mupa888Short for isl_multi_union_pw_aff!. object.
The SET_PARALLEL transformation simply insert a “mark” node, with an isl_id name "parallel" and the code generatation handles the rest.
Similarly to SET_PARALLEL, SET_LOOP_OPT is not strictly speaking a (polyhedral) transformation. It is implemented by simply calling the corresponding ISL function, and influences how code ISL generates code.
4.5. Legality Checking
Flow of legality check for two examples.
After a transformation concludes, the legality check is invoked on the same node. The legality check implements the principles described in Section 2.2. Each step of the legality check in Algorithm 1 can be implemented with a single ISL function. Algorithm 1 is applied after every transformation with the exception of SET_PARALLEL. For the legality of parallel loops the entry in corresponding to the loop should be , and this can be achieved with a slight modification to Algorithm 1: we use the partial schedule instead of the complete schedule in steps 1 and 2, and in step 4 we compute and and check if they are both empty in step 5. Figure 4 shows two examples of the process of checking legality. Checking legality starts with the dependency graph at the top, which, in case of Listing 1, is one vertex with a loop edge.
The domain and range of the map (below ), which indicates that statement instance depends on , are transformed according to the schedule .
The resulting transformed map is shown below (dotted rectangle).
If the difference between the (transformed) domain and the (transformed) range, denoted by , is lexicographicaly positive, the schedule is legal (left side, the from Figure 2(c)), otherwise the schedule is not legal (right side, the from Figure 2(d)).
4.6. Code Generation
Code generation is based on the pet_loopback.c file from the PET repository, which extracts the polyhedral representation from an input C file, and generates the “same” code using the original schedule.
Apart from the obvious tweaking to use the modified schedule instead of the original one, and to make it cooperate with the bookkeeping setup we use in Tadashi’s back-end, the SET_PARALLEL transformation required additional modifications to the core code generation.
5. Evaluation
We evaluate the capabilities of Tadashi by verifying functional correctnes, using two different ML solutions on top of Tadashi to showcase how Tadashi enables ML solutions to speedup different applications and benchmark suites. Then we demonstrate how Tadashi can be integrated with existing scientific computing codes, illustrating Tadashi’s potential to enabling AI-based automated code generation with guaranteed correctnesss for scientific applications, and finally measure the overhead.
5.1. Empirical Functional Testing
We provided correctness guarantees of the polyhedral method we developed for Tadashi in Section 4.5. However, we discuss in this section how we empirically do functional testing to assure the correct implementation of Tadashi. We evaluated Tadashi by applying a range of primitive (single) and composite (combined) transformations on kernels from the Polybench (Reisinger, 2024) benchmark suite.
To verify the correctness of implementation of Tadashi,
first, all Polybench benchmarks were compiled and executed with the small dataset999Using -DSMALL_DATASET!. and with the output saved.101010Using -DPOLYBENCH_DUMP_ARRAYS!. In an automated manner, we generate 3 random transformations,
rollback the illegal ones, compile, run, and compare the output of the transformed program to the original.
Post-inspecting confirmed the outputs to be identical.
5.2. Using Tadashi with ML Solutions
To demonstrate the practical utility of Tadashi in real-world ML workflows, we developed and deployed three illustrative examples using our framework. The first is a simple proof-of-concept implementation of an optimal transformation search (POC-search), designed to highlight Tadashi’s ability to apply, verify and evaluate transformations: While experimenting with transformations, we noticed fission and tiling (2D and 3D) can often speed up Polybench kernels,111111Using the extra large dataset. so with a fixed tile size, we implemented a brute-force search, applying fission and tiling when ever possible, rolling back and discarding illegal results to obtain speedups presented in Figure 5. The atax, bicg, gesummv, and jacobi-1d kernels are not included since their execution time is under 0.01s and, as such, improvements on the results can be caused by noise in the CPU. A speedup of 1x indicates that no valid operations were found (cholesky, deriche, durbin, ludcmp, seidel-2d) or that the transformation did not result in a speedup (gesummv).
POC algorithm speedups
The other two are representative ML models that showcase how Tadashi enables verifiably correct code generation in the context of ML. Each of these examples illustrates different aspects of Tadashi’s design and capabilities.
5.2.1. Monte-Carlo Tree Search
We have adopted Monte Carlo Tree Search (MCTS) (Coulom, 2007) as a baseline method for discovering effective transformations and compositions of transformations. MCTS algorithm iteratively builds a search tree where nodes represent states of the code after applying specific transformations (such as tiling, fusion, or interchange), and edges represent the transformations themselves. More precisely, we break each transformation down into three steps: selecting which node to transform, which transformation to apply and which parameters (e.g., tile size) to select for a given transformation. This approach naturally handles the composition of transformations: applicability and effect of each transformation depends on previously applied ones. MCTS requires no domain-specific heuristics beyond a performance evaluation function, thus allowing it to discover non-intuitive transformations.
Furthermore we employ the Upper Confidence Bound for Trees (UCT) (Kocsis and Szepesvári, 2006) strategy to balance exploration of novel transformation sequences with exploitation of already discovered effective patterns.
Figure 6 illustrates the pace of discovery of effective transformations for a range of kernels from the Polybench suite. We excluded the loop parallelization transformation since in most cases it is easy to do manually, thus we wanted to focus on single-core performance improvement. Each chart is a single run of Tadashi on an AMD EPYC 9684X 96-Core Processor. The main cost of running Tadashi is the actual execution of the transformed code. We aimed at execution time of about 10 seconds for a given kernel, to strike a balance between reliability of measurements and reasonable turn-around time. Since the execution time can be set arbitrary, we report performance improvement over the number of evaluation steps.
Plot of speedups discovered by MCTS on different benchmarks
In line with our intuition for the POC approach, MCTS quickly discovered that as stand-alone transformations, interchange (e.g., for gemm) and tiling (e.g., for floyd-warshall) are often very effective. In terms of compositions, we see interchange or full split combined with tiling being discovered on multiple occasions. As a concrete example, transformations resulting in 6x performance improvement in trmm kernel were discovered as follows:
trs = [ [3, "FULL_SPLIT"], [1, "TILE1D", 3], [5, "INTERCHANGE"] ]
The first number is the index of a node in the polyhedral schedule tree, last number indicates the transformation parameter. More examples can be found in the examples/evaluation/data directory in the project repository.
5.2.2. Evolutionary Computation
In addition to the MCTS, we explored using a Genetic Programming (GP)-based (Vanneschi and Silva, 2023) approach to search for a good set of transformations. Unlike MCTS, GP is allowed to use all transformations provided by Tadashi. We limit the number of iterations to 20 and use a population size of 20, i.e., the algorithm does a total of 400 evaluations to obtain the final list of operations.
In this implementation, a model is represented as a list of operations to be applied to a SCoP. The models are initialized as an empty list of transformations. During training, the models can be mutated to append a randomly selected transformation to their list or lose one operation, and can also be combined with other models by picking a cutting point from both models’ lists and swapping the latter halves, creating two new models. Any invalid models generated this way are discarded.
The GP experiments are still in a preliminary stage, and while we only obtained results for two of the Polybench benchmarks (gemm and jacobi-2d), the results are promising. On gemm, GP obtained a 18.2x speedup (4.0x, without using the parallel transformation) after 12 iterations (240 evaluations). On jacobi-2d, GP obtained a 2.8x speedup after 8 iterations (160 evaluations). These results indicate that GP can beat the results previously seen in Figure 5.
5.2.3. Transformations Found by Tadashi
The following lists show two examples of the Tadashi transformations that GP found to speedup gemm and jacobi-2d by 18.2x and 2.8x, respectively. Additional details of transformations that Tadashi discover are included in the supplementary material, as permitted for accepted papers. Each row contains the node index, the transformation and its arguments.
gemm = [ [1, "FULL_SHIFT_PARAM", 2, 16], [2, "FULL_FUSE", ], [1, "SET_PARALLEL", 6], [2, "FULL_SHIFT_PARAM", 1, 48], [2, "TILE2D", 32, 32], [3, "SET_LOOP_OPT", 0, 3], ] jacobi-2d = [ [9, "TILE1D", 32], [4, "TILE2D", 32, 32], [10, "TILE2D", 32, 32], [10, "TILE1D", 32], [13, "INTERCHANGE"], ]
5.3. Integration with Scientific Applications
To evaluate the applicability of Tadashi into more realistic scientific computing scenarios, beyond simple benchmarks overspecialized for polyhedral optimization, we integrated it with two representative applications: miniAMR (Vaughan and Barrett, 2015) and FDTD electromagnetic simulation from (Sullivan, 2013). These codes, while manageable in size, reflect the complexity and structure of production-level scientific software, providing valuable insight into how Tadashi can support correctness and agency in practical settings.
miniAMR is a proxy application developed as part of the Mantevo project, designed to emulate the behavior of Adaptive Mesh Refinement (AMR) codes used in large-scale physics simulations. It features block-structured grids, dynamic refinement and coarsening, and stencil-based updates—characteristics commonly found in production AMR codes.
The FDTD electromagnetic simulation uses the Finite-Difference Time-Domain (FDTD) method. It implements a full 3D simulation kernel including absorbing boundary conditions, and serves as a self-contained, practical example of a numerical method widely used in computational electromagnetics.
The FDTD electromagnetic simulation is relatively simple and self-contained, with a clear computational kernel, which made it well suited for Tadashi-driven optimization–resulting in a 13x speedup. miniAMR, on the other hand, is significantly more complex, featuring dynamic mesh refinement and distributed data structures. Despite these challenges, Tadashi was able to produce a 1.05x (5%) speedup, demonstrating utility even in more intricate scientific codes.
The modest 1.05x speedup in miniAMR highlights the need for more sophisticated machine learning techniques and a more elaborate setup to fully exploit the potential of Tadashi in complex applications. Specifically, a more refined strategy for distributing the evaluations and measuring the runtime across multiple nodes would likely yield better scaling of ML methods. Fortunately, Tadashi fully supports such a distributed setup, leveraging tools like MPI4Py.futures to enable parallel execution and scalable optimizations across a cluster. This capability will be crucial as we continue to refine our approach and tackle even more complex scientific workloads.
5.4. Overhead
Breakdown of the results
Throughput
ML training algorithms need a vast number of code transformations, including the legality check and code generation, it would be informative to see how fast Tadashi performs these functions. With this in mind, we devise two experiments to measure a typical training cycle: one which involves a single polyhedral transformation and another with a sequence of 10 transformations (and legality checks), with a code generation at the end. A typical training cycle would cover functions performed by Tadashi, such as, extracting the polyhedral representation of the SCoPs, performing the transformation, checking the legality, generating code, as well as, compiling and running the program; steps with execution times which neither the user nor Tadashi can influence or circumvent. For both experiments, we use Polybench with the default dataset size, based on the reasoning that a smaller dataset would fit into cache memory and shrink the runtime to an extent where optimization would not be required.
The breakdown of the results can be found in Figure 7. The dominance of the runtime of the programs is obvious in more then half of the apps. As for the other benchmarks, with miniscule runtimes, we found that the functions performed by Tadashi always occupied less then half of the breakdown. Based on these numbers, using Tadashi should not be considerably more prohibitive than any other workflow which involves compiling and/or running the program.
The overhead measurements are done on a regular desktop with a 12 core AMD Ryzen 5 5600X 6-Core Processor and 64GiB of memory, running Arch Linux, with python 3.12.5-1, libisl 0.27-1, llvm 18.1.8-4 system packages, and Tadashi’s CMake script using the current master branch of PET121212From https://repo.or.cz/pet.git. tagged as pet 0.11.8.
To provide a better insight into Tadashi’s performance, we also calculate throughput from the breakdown data, which is shown in Figure 8. Only the number of transformations and legality checks are counted, since it can happen more often than SCoP extraction and code generation, and the latter two would also be accompanied by compilation and program execution, which, as we saw in Figure 7, would dominate the runtime. The results show Tadashi capability of sustaining 100s of iterations per second, which would be very practical for production ML solutions where 100,000s 100,000,000s of training examples or direct evaluations in RL would be the norm.
6. Conclusion
In this work we aim to lower the barrier for using ML in code auto generation, with guarantees on correctness. We rely on the polyhedral methods to generate code based on compositional loop transformations with mathematically proven guarantee of functional correctness. To this end we introduce Tadashi, a library which exposes an API simplified to the polyhedral methods, targeting the ML engineers and researchers not familiar with it. By using Tadashi, ML practitioners can design ML-driven methods based on the capability of Tadashi in sampling of the space of legal loop transformations.
While our initial experiments with machine learning methods validated the utility of Tadashi and the approach in general, they also revealed the vastness of the search space: some kernels in real applications translated into hundreds of polyhedral schedule tree nodes, each of which could be modified by a combination of numerous transformations and parameters. This realization, in turn, necessitates future work to create a solution that would allow Tadashi to run at the scale of large HPC clusters, enabling the discovery of polyhedral optimizations that go beyond those described in the literature. We leave this exploration for future work.
7. Acknowledgement
ChatGPT (OpenAI, [n. d.]) was used for Grammar and Spelling correction.
Acknowledgements.
We are eternally grateful to Sven (skimo) Verdoolaege for his invaluable help by replying to our (often silly) questions on the ISL mailing list.References
- (1)
- exa ([n. d.]) [n. d.]. The Exascale Software Portfolio | Science & Technology Review. https://str.llnl.gov/past-issues/february-2021/exascale-software-portfolio
- 202 (2025) 2025. Sakana AI’s Claim That It Can ’make Things 100 Times Faster’ Has Been Verified Online and It Has Been Pointed out That It Will Actually ’Make Things 3 Times Slower’. http://gigazine.net/gsc_news/en/20250225-sakana-ai-cuda-engineer-walks-back/
- Aho et al. (2007) Alfred V. Aho, Monica S. Lam, Ravi Sethi, and Jeffrey D. Ullman. 2007. Compilers: Principles, Techniques, & Tools (second edition ed.). Pearson/Addison Wesley, Boston. http://catdir.loc.gov/catdir/toc/ecip0618/2006024333.html
- Allamanis et al. (2024) Miltiadis Allamanis, Sheena Panthaplackel, and Pengcheng Yin. 2024. Unsupervised Evaluation of Code LLMs with Round-Trip Correctness. In ICML. OpenReview.net.
- Ashouri et al. (2018) Amir H. Ashouri, William Killian, John Cavazos, Gianluca Palermo, and Cristina Silvano. 2018. A Survey on Compiler Autotuning using Machine Learning. ACM Comput. Surv. 51, 5, Article 96 (sep 2018), 42 pages. doi:10.1145/3197978
- Baghdadi et al. (2021) Riyadh Baghdadi, Massinissa Merouani, Mohamed-Hicham Leghettas, Kamel Abdous, Taha Arbaoui, Karima Benatchba, and Saman P. Amarasinghe. 2021. A Deep Learning Based Cost Model for Automatic Code Optimization. In Proceedings of the Fourth Conference on Machine Learning and Systems, MLSys 2021, Virtual, April 5-9, 2021, Alex Smola, Alex Dimakis, and Ion Stoica (Eds.). mlsys.org. https://proceedings.mlsys.org/paper_files/paper/2021/hash/d9387b6d643efb25132be36f7b908d96-Abstract.html
- Beer et al. (2024) Robin Beer, Alexander Feix, Tim Guttzeit, Tamara Muras, Vincent Müller, Maurice Rauscher, Florian Schäffler, and Welf Löwe. 2024. Examination of Code generated by Large Language Models. arXiv:2408.16601 [cs.SE] https://arxiv.org/abs/2408.16601
- Ben-Nun et al. (2018) Tal Ben-Nun, Alice Shoshana Jakobovits, and Torsten Hoefler. 2018. Neural Code Comprehension: A Learnable Representation of Code Semantics. In Proceedings of the 32nd International Conference on Neural Information Processing Systems (NIPS’18). Curran Associates Inc., Red Hook, NY, USA, 3589–3601.
- Bondhugula et al. (2008) Uday Bondhugula, Albert Hartono, J Ramanujam, and P Sadayappan. 2008. A Practical Automatic Polyhedral Parallelizer and Locality Optimizer. (2008), 13.
- Cai et al. (2023) Zonghui Cai, Xiao Yang, MengChu Zhou, Zhi-Hui Zhan, and Shangce Gao. 2023. Toward explicit control between exploration and exploitation in evolutionary algorithms: A case study of differential evolution. Information Sciences 649 (2023), 119656. doi:10.1016/j.ins.2023.119656
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, Alex Ray, Raul Puri, Gretchen Krueger, Michael Petrov, Heidy Khlaaf, Girish Sastry, Pamela Mishkin, Brooke Chan, Scott Gray, Nick Ryder, Mikhail Pavlov, Alethea Power, Lukasz Kaiser, Mohammad Bavarian, Clemens Winter, Philippe Tillet, Felipe Petroski Such, Dave Cummings, Matthias Plappert, Fotios Chantzis, Elizabeth Barnes, Ariel Herbert-Voss, William Hebgen Guss, Alex Nichol, Alex Paino, Nikolas Tezak, Jie Tang, Igor Babuschkin, Suchir Balaji, Shantanu Jain, William Saunders, Christopher Hesse, Andrew N. Carr, Jan Leike, Josh Achiam, Vedant Misra, Evan Morikawa, Alec Radford, Matthew Knight, Miles Brundage, Mira Murati, Katie Mayer, Peter Welinder, Bob McGrew, Dario Amodei, Sam McCandlish, Ilya Sutskever, and Wojciech Zaremba. 2021. Evaluating Large Language Models Trained on Code. arXiv (2021). arXiv:2107.03374 [cs.LG]
- Coello (2022) Carlos A. Coello Coello. 2022. Constraint-handling techniques used with evolutionary algorithms. In Proceedings of the Genetic and Evolutionary Computation Conference Companion (Boston, Massachusetts) (GECCO ’22). Association for Computing Machinery, New York, NY, USA, 1310–1333. doi:10.1145/3520304.3533640
- Cotroneo et al. (2024) Domenico Cotroneo, Alessio Foggia, Cristina Improta, Pietro Liguori, and Roberto Natella. 2024. Automating the correctness assessment of AI-generated code for security contexts. Journal of Systems and Software 216 (2024), 112113. doi:10.1016/j.jss.2024.112113
- Coulom (2007) Rémi Coulom. 2007. Efficient Selectivity and Backup Operators in Monte-Carlo Tree Search. In Computers and Games, H. Jaap van den Herik, Paolo Ciancarini, and H. H. L. M. (Jeroen) Donkers (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 72–83.
- Cummins et al. (2020) Chris Cummins, Zacharias V. Fisches, Tal Ben-Nun, Torsten Hoefler, and Hugh Leather. 2020. ProGraML: Graph-based Deep Learning for Program Optimization and Analysis. (March 2020). doi:10.48550/arXiv.2003.10536
- Cummins et al. (2023) Chris Cummins, Volker Seeker, Dejan Grubisic, Mostafa Elhoushi, Youwei Liang, Baptiste Roziere, Jonas Gehring, Fabian Gloeckle, Kim Hazelwood, Gabriel Synnaeve, and Hugh Leather. 2023. Large Language Models for Compiler Optimization. arXiv:2309.07062 [cs.PL]
- Cummins et al. (2024) Chris Cummins, Volker Seeker, Dejan Grubisic, Baptiste Roziere, Jonas Gehring, Gabriel Synnaeve, and Hugh Leather. 2024. Meta Large Language Model Compiler: Foundation Models of Compiler Optimization. doi:10.48550/arXiv.2407.02524 arXiv:2407.02524 [cs]
- Dehaerne et al. (2022) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. doi:10.1109/ACCESS.2022.3196347
- Du et al. (2024) Xueying Du, Mingwei Liu, Kaixin Wang, Hanlin Wang, Junwei Liu, Yixuan Chen, Jiayi Feng, Chaofeng Sha, Xin Peng, and Yiling Lou. 2024. Evaluating Large Language Models in Class-Level Code Generation. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 81, 13 pages. doi:10.1145/3597503.3639219
- Feautrier (1991) Paul Feautrier. 1991. Dataflow Analysis of Array and Scalar References. (1991), 37. Introduction of SCoP.
- Feautrier (1996) Paul Feautrier. 1996. Some Efficient Solutions to the Affine Scheduling Problem Part I One-dimensional Time. International Journal of Parallel Programming 21 (Aug. 1996). doi:10.1007/BF01407835
- Feautrier (1997) Paul Feautrier. 1997. Some Efficient Solutions to the Affine Scheduling Problem Part II Multidimensional Time. International Journal of Parallel Programming 21 (Jan. 1997). doi:10.1007/BF01379404
- Github (2022) Github. 2022. GitHub Copilot Your AI Pair Programmer. https://github.com/features/copilot
- Grosser et al. (2012) Tobias Grosser, Armin Groesslinger, and Christian Lengauer. 2012. Polly - Performing Polyhedral Optimizations on a Low-Level Intermediate Representation. Parallel Process. Lett. 22, 04 (Dec. 2012), 1250010. doi:10.1142/S0129626412500107
- Heroux (2024) Michael A. Heroux. 2024. Scalable Delivery of Scalable Libraries and Tools: How ECP Delivered a Software Ecosystem for Exascale and Beyond. Computing in Science & Engineering 26, 1 (Jan. 2024), 9–18. doi:10.1109/MCSE.2024.3384937
- Jiang et al. (2024) Juyong Jiang, Fan Wang, Jiasi Shen, Sungju Kim, and Sunghun Kim. 2024. A Survey on Large Language Models for Code Generation. arXiv:2406.00515 [cs.CL]
- Kocsis and Szepesvári (2006) Levente Kocsis and Csaba Szepesvári. 2006. Bandit based monte-carlo planning. In Proceedings of the 17th European Conference on Machine Learning (Berlin, Germany) (ECML’06). Springer-Verlag, Berlin, Heidelberg, 282–293. doi:10.1007/11871842_29
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Proceedings of the 34th International Conference on Neural Information Processing Systems (NIPS ’20). Curran Associates Inc., Red Hook, NY, USA, 9459–9474.
- Liang et al. (2023) Youwei Liang, Kevin Stone, Ali Shameli, Chris Cummins, Mostafa Elhoushi, Jiadong Guo, Benoit Steiner, Xiaomeng Yang, Pengtao Xie, Hugh Leather, and Yuandong Tian. 2023. Learning compiler pass orders using coreset and normalized value prediction. In Proceedings of the 40th International Conference on Machine Learning (Honolulu, Hawaii, USA) (ICML’23). JMLR.org, Article 855, 17 pages.
- Mankowitz et al. (2023) Daniel J. Mankowitz, Andrea Michi, Anton Zhernov, Marco Gelmi, Marco Selvi, Cosmin Paduraru, Edouard Leurent, Shariq Iqbal, Jean-Baptiste Lespiau, Alex Ahern, Thomas Köppe, Kevin Millikin, Stephen Gaffney, Sophie Elster, Jackson Broshear, Chris Gamble, Kieran Milan, Robert Tung, Minjae Hwang, Taylan Cemgil, Mohammadamin Barekatain, Yujia Li, Amol Mandhane, Thomas Hubert, Julian Schrittwieser, Demis Hassabis, Pushmeet Kohli, Martin Riedmiller, Oriol Vinyals, and David Silver. 2023. Faster Sorting Algorithms Discovered Using Deep Reinforcement Learning. Nature 618, 7964 (June 2023), 257–263. doi:10.1038/s41586-023-06004-9
- Munoz et al. (2024) Michael Munoz, Juan Torre, Stefany Lopez, Stiven Herrera, and Christian Uribe. 2024. Comparative Study of AI Code Generation Tools: Quality Assessment and Performance Analysis. LatIA 2 (07 2024), 104. doi:10.62486/latia2024104
- Mustafa (2022) Dheya Mustafa. 2022. A Survey of Performance Tuning Techniques and Tools for Parallel Applications. IEEE Access 10 (2022), 15036–15055. doi:10.1109/ACCESS.2022.3147846
- OpenAI ([n. d.]) OpenAI. [n. d.]. ChatGPT. https://chat.openai.com
- Ouyang et al. (2024) Anne Ouyang, Simon Guo, and Azalia Mirhoseini. 2024. KernelBench: Can LLMs Write GPU Kernels? In None. https://scalingintelligence.stanford.edu/blogs/kernelbench/
- Phan (2019) Hung Phan. 2019. Self Learning from Large Scale Code Corpus to Infer Structure of Method Invocations. https://arxiv.org/abs/1909.03147v1
- Pouchet et al. (2011) Louis-Noël Pouchet, Uday Bondhugula, Cédric Bastoul, Albert Cohen, J. Ramanujam, P. Sadayappan, and Nicolas Vasilache. 2011. Loop transformations: convexity, pruning and optimization. In Proceedings of the 38th Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages (Austin, Texas, USA) (POPL ’11). Association for Computing Machinery, New York, NY, USA, 549–562. doi:10.1145/1926385.1926449
- Reisinger (2024) Matthias Reisinger. 2024. MatthiasJReisinger/PolyBenchC-4.2.1. https://github.com/MatthiasJReisinger/PolyBenchC-4.2.1
- Rogowski et al. (2023) Marcin Rogowski, Samar Aseeri, David Keyes, and Lisandro Dalcin. 2023. Mpi4py.Futures: MPI-Based Asynchronous Task Execution for Python. IEEE Transactions on Parallel and Distributed Systems 34, 2 (Feb. 2023), 611–622. doi:10.1109/TPDS.2022.3225481
- Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessi, Roberta Raileanu, Maria Lomeli, Eric Hambro, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language Models Can Teach Themselves to Use Tools. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 68539–68551. https://proceedings.neurips.cc/paper_files/paper/2023/file/d842425e4bf79ba039352da0f658a906-Paper-Conference.pdf
- Shido et al. (2019) Yusuke Shido, Yasuaki Kobayashi, Akihiro Yamamoto, Atsushi Miyamoto, and Tadayuki Matsumura. 2019. Automatic Source Code Summarization with Extended Tree-LSTM. In 2019 International Joint Conference on Neural Networks (IJCNN). 1–8. doi:10.1109/IJCNN.2019.8851751
- Silver et al. (2016) David Silver, Aja Huang, Chris J. Maddison, Arthur Guez, Laurent Sifre, George van den Driessche, Julian Schrittwieser, Ioannis Antonoglou, Veda Panneershelvam, Marc Lanctot, Sander Dieleman, Dominik Grewe, John Nham, Nal Kalchbrenner, Ilya Sutskever, Timothy Lillicrap, Madeleine Leach, Koray Kavukcuoglu, Thore Graepel, and Demis Hassabis. 2016. Mastering the game of Go with deep neural networks and tree search. Nature 529, 7587 (01 Jan 2016), 484–489. doi:10.1038/nature16961
- Stein et al. (2020) Aviel J. Stein, Levi Kapllani, Spiros Mancoridis, and Rachel Greenstadt. 2020. Exploring Paraphrasing Techniques on Formal Language for Generating Semantics Preserving Source Code Transformations. In 2020 IEEE 14th International Conference on Semantic Computing (ICSC). 242–248. doi:10.1109/ICSC.2020.00051
- Sullivan (2013) Dennis Michael Sullivan. 2013. Electromagnetic Simulation Using the FDTD Method (second edition ed.). IEEE Press : John Wiley & Sons, Inc, Hoboken, New Jersey.
- Sutton and Barto (2018) Richard S Sutton and Andrew G Barto. 2018. Reinforcement Learning: An Introduction. MIT press.
- Trümper et al. (2023) Lukas Trümper, Tal Ben-Nun, Philipp Schaad, Alexandru Calotoiu, and Torsten Hoefler. 2023. Performance Embeddings: A Similarity-Based Transfer Tuning Approach to Performance Optimization. In Proceedings of the 37th ACM International Conference on Supercomputing (Orlando, FL, USA) (ICS ’23). Association for Computing Machinery, New York, NY, USA, 50–62. doi:10.1145/3577193.3593714
- Vanneschi and Silva (2023) Leonardo Vanneschi and Sara Silva. 2023. Genetic Programming. Springer International Publishing, Cham, 205–257. doi:10.1007/978-3-031-17922-8_8
- Vaughan and Barrett (2015) Courtenay T. Vaughan and Richard F. Barrett. 2015. Enabling Tractable Exploration of the Performance of Adaptive Mesh Refinement. In 2015 IEEE International Conference on Cluster Computing. 746–752. doi:10.1109/CLUSTER.2015.129
- Verdoolaege (2010) Sven Verdoolaege. 2010. Isl: An Integer Set Library for the Polyhedral Model. In Mathematical Software – ICMS 2010, Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama (Eds.). Springer, Berlin, Heidelberg, 299–302. doi:10.1007/978-3-642-15582-6_49
- Verdoolaege et al. (2013) Sven Verdoolaege, Juan Carlos Juega, Albert Cohen, José Ignacio Gómez, Christian Tenllado, and Francky Catthoor. 2013. Polyhedral Parallel Code Generation for CUDA. ACM Trans. Archit. Code Optim. 9, 4 (Jan. 2013), 54:1–54:23. doi:10.1145/2400682.2400713
- Verdoolaege and Grosser (2012) Sven Verdoolaege and Tobias Grosser. 2012. Polyhedral Extraction Tool. (2012), 8.
- Verdoolaege et al. (2014) Sven Verdoolaege, Serge Guelton, Tobias Grosser, and Albert Cohen. 2014. Schedule Trees. In International Workshop on Polyhedral Compilation Techniques. Vienna, Austria. https://lirias.kuleuven.be/1656334
- Verdoolaege et al. (2021) Sven Verdoolaege, Polly Labs, and KU Leuven. 2021. Presburger Formulas and Polyhedral Compilation. (2021), 174.
- Wang et al. (2024) Ruotong Wang, Ruijia Cheng, Denae Ford, and Thomas Zimmermann. 2024. Investigating and Designing for Trust in AI-powered Code Generation Tools. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (Rio de Janeiro, Brazil) (FAccT ’24). Association for Computing Machinery, New York, NY, USA, 1475–1493. doi:10.1145/3630106.3658984
- Wei et al. (2024) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2024. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
- While and Hingston (2013) Lyndon While and Philip Hingston. 2013. Usefulness of infeasible solutions in evolutionary search: An empirical and mathematical study. In 2013 IEEE Congress on Evolutionary Computation. 1363–1370. doi:10.1109/CEC.2013.6557723
- Zhou et al. (2023) Shuyan Zhou, Uri Alon, Frank F. Xu, Zhengbao Jiang, and Graham Neubig. 2023. DocPrompting: Generating Code by Retrieving the Docs. In ICLR. OpenReview.net.